### v1

# Partially Retargeted Balancing Weights for Causal Effect Estimation Under Positivity Violations

Abstract
Positivity violations pose significant challenges for causal effect estimation with observational data. Under positivity violations, available methods result in either treatment effect estimators with substantial statistical bias and variance or estimators corresponding to a modified estimand and target population that is misaligned with the original research question. To address these challenges, we propose partially retargeted balancing weights, which yield reduced estimator variance under positivity violations by modifying the target population for only a subset of covariates. Our weights can be derived under a novel relaxed positivity assumption allowing the calculation of valid balancing weights even when positivity does not hold. Our proposed weighted estimator is consistent for the original target estimand when either 1) the implied propensity score model is correct; or 2) the subset of covariates whose population is not modified contains all treatment effect modifiers. When these conditions do not hold, our estimator is consistent for a slightly modified treatment effect estimand. Furthermore, our proposed weighted estimator has reduced asymptotic variance when positivity does not hold. We evaluate our weights and corresponding estimator through applications to synthetic data, an EHR study, and when transporting an RCT treatment effect to a Midwestern population.
Keywords: causal inference, direct balancing, overlap, transportability, observational studies
* barna126@umn.edu
Introduction
Confounding is a key challenge when estimating causal effects of binary treatments with observational data. While many methods exist for controlling confounding by balancing covariates between treatment groups, these methods do not fully control confounding when there are positivity violations. Positivity violations occur when there exists at least one combination of covariate values which perfectly predicts treatment assignment, a phenomenon which is especially likely to occur with high-dimensional covariates (D’Amour and Franks,, 2021). When there are positivity violations, inverse probability weighting (IPW) estimators (Rosenbaum and Rubin,, 1983; Hahn,, 1998; Robins et al.,, 2000; Hirano and Imbens,, 2001; Hirano et al.,, 2003; Imbens,, 2004) have substantial statistical bias and inflated variance due to extreme weights. A proposed alternative to IPW, direct balancing weights, (Hainmueller,, 2012; Zubizarreta,, 2015; Wang and Zubizarreta,, 2020) tend to yield reduced error and variance compared to IPW methods through achieving exact covariate balance for a prespecified set of covariates. However, when there are positivity violations there is often no solution to the direct balancing optimization problem prohibiting their use altogether.
To mitigate the impacts of positivity violations on causal effect estimation, a variety of methods, such as propensity score trimming, matching weights, and overlap weights, have been developed that modify the estimand to one whose target population has increased overlap to achieve estimators with reduced variance (Crump et al.,, 2009; Li and Greene,, 2013; Yang and Ding,, 2018; Li et al.,, 2018, 2019). In fact, overlap weights (Li et al.,, 2018) do not require positivity at all by targeting the so-called ‘overlap population’; this result in an estimator with minimum asymptotic variance among weighted treatment effects under homoscedasticity. However, these methods target a population that is often substantially different than the original population of interest such that the corresponding treatment effect estimate may be misaligned with the original research question. Yet, a key benefit of causal effect estimation with observational data is the ability to evaluate causal effects in a meaningful population of interest (e.g., transporting effects) such that modifying the estimand is either undesirable or challenging to interpret. Furthermore, the corresponding estimators are biased for the original estimand when there is treatment effect heterogeneity. Modified treatment policy estimands result in improved estimator performance by defining an entirely different estimand arrived at by imagining modifications to treatment (Kennedy,, 2019). While these estimators tend to perform well, these modified treatments tend to be more challenging to conceptualize than binary treatments.
Thus, under positivity violations, methods either 1) target an estimand of interest but result in poor estimator performance or technical issues (e.g., balancing weights); or 2) alleviate these technical issues by targeting a modified estimand that may be misaligned with the original estimand of interest (e.g., overlap weights). To address these challenges of estimator variance, bias, and interpretation, we propose a novel direct balancing approach whose optimization problem yields a solution under positivity violations by modifying the target population for only a subset of the covariates. In doing so, our proposed weights achieve exact covariate balance between treated and control groups for all covariates; in contrast, minimal weights (Wang and Zubizarreta,, 2020) achieve a solution through by focusing on approximate, rather than exact, covariate balance. The target population under our approach can be modified as minimally as possible to achieve a solution and the corresponding estimator tends to minimize bias and variance with respect to the modified estimand. When the target population is instead modified for all covariates, overlap weights are a special case of our proposed weights. Our proposed estimator is consistent for the original estimand of interest when the either the implied propensity score model is correct or the subset of covariates whose population is modified are not treatment effect modifiers. We derive asymptotic results which show that 1) our proposed approach explicitly relaxes the positivity assumption; and 2) our proposed estimator has reduced asymptotic variance when there are positivity violations. We propose a design- and model-based procedure for selecting the subset of the covariates not balanced to the target population and we evaluate the corresponding estimators through applications to synthetic data, the MIMIC-III study, and when transporting the effect of a healthcare hot spotting study.
The remainder of the paper is organized as follows. In Section 2, we introduce notation and methodological background on direct balancing and minimal weights. We introduce our proposed partially retargeted direct balancing approach and explore the corresponding relaxed positivity assumption, interpretation of the implied propensity score, and outcome models in Section 3. In Section 4, we discuss the asymptotic properties of our proposed estimator and in Section 5 we discuss adaptations of proposed method to other estimation problems. We discuss implementation details and propose design- and model-based estimators for our method in Section 6. In Sections 7, 8, and 9, we apply our methods to synthetic data, an EHR study, and transporting a health care hotspotting RCT effect.
Methodological background
Notation, assumptions, and covariate balance
Consider an independent sample $\{(Z_{i},Y_{i},\bm{X}_{i})\}_{i=1}^{n}$ of size $n$ from a population where $Z_{i}=z$, indicates belonging to the treatment ($z=1$) or control group ($z=0$), $Y_{i}$ is the outcome of interest, and $\bm{X}_{i}=(X_{i1},\ldots X_{ip})$ is a $p$ length vector of pre-treatment covariates. Let $Y(z)$ be the potential outcome that would be observed if assigned to treatment group $z$ (Neyman, (1990); Rubin, (1974, 1978); Hernan and Robins, (2024)). We assume the standard stable unit treatment value assumption (SUTVA) which implies that $Y_{i}=Y_{i}(Z_{i})$ such that we only observe one potential outcome for each individual. We focus on the average treatment effect (ATE), $\tau=E[Y(1)-Y(0)]$ as the estimand of interest, although this work can be readily extended to other estimands of interest which we discuss in Section 5. Let $e(\bm{x})=\Pr(Z_{i}=1|\bm{X}_{i}=\bm{x})$ be the propensity score. For the identification of $\tau$ the following assumptions are typically made:
Assumption 1: Strong ignorability.
$\{Y(0),Y(1)\}{\,\perp\!\!\!\perp\,}Z|\bm{X}$ and
Assumption 2: Positivity.
$0<e(\bm{x})<1$ for all $\bm{x}$.
The focus of our work surrounds the positivity assumption, specifically, scenarios where this assumption does not hold. Let $\mu_{z}(\bm{x})=E[Y(z)|\bm{X}=\bm{x}]$ and $\tau(\bm{x})=\mu_{1}(\bm{x})-\mu_{0}(\bm{x})$. Then the following identification holds with the above assumptions:
$\tau$ $=E\left\{\frac{\mu_{1}(\bm{X})Z}{e(\bm{X})}-\frac{\mu_{0}(\bm{X})(1-Z)}{1-e(\bm{X})}\right\}=E\left\{\frac{Y(1)Z}{e(\bm{X})}-\frac{Y(0)(1-Z)}{1-e(\bm{X})}\right\}=E\left\{\frac{YZ}{e(\bm{X})}-\frac{Y(1-Z)}{1-e(\bm{X})}\right\},$
where the sample version of the last equation corresponds to the standard inverse probability weighting (IPW) estimator of the ATE. For a treatment effect estimator be unbiased, the expectations $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ need to be balanced between the treated and control populations and to the target population. IPW balance all functions of $\bm{x}$ in this way such that the following equations hold,
$E\left\{\frac{Zb(\bm{X})}{e(\bm{X})}\right\}=E\{b(\bm{X})\}\>\>$ $\text{and}\>\>E\left\{\frac{(1-Z)b(\bm{X})}{1-e(\bm{X})}\right\}=E\{b(\bm{X})\}$ (1)
$E\left\{\frac{Zb(\bm{X})}{e(\bm{X})}\right\}$ $=E\left\{\frac{(1-Z)b(\bm{X})}{1-e(\bm{X})}\right\}$ (2)
for any function $b(\bm{x})$, which we reference as three-way balance (Chan et al.,, 2016). Thus, IPW satisfy three-way balance of both $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$, regardless of their true forms. Specifically, Equation (1) balances the weighted treated and control group expectations of $b(\bm{x})$ to the population expectation of $b(\bm{x})$ and Equation (2) balances the weighted treated and control group expectations of $b(\bm{x})$ to each other. Thus, Equation (1) and (2) serve different purposes; while Equation (1) implies Equation (2), Equation (1) primarily serves to balance $b(\bm{x})$ in the treated and control groups to $b(\bm{x})$ to the estimand target population. In contrast, Equation (2) only ensures that $b(\bm{x})$ is balanced between treated and control groups. When only Equation (2) holds, the corresponding treatment effect estimator will be consistent for a modified target population and causal estimand but could not be consistent for the ATE.
However, when there is a lack of overlap, even approximate balance of the empirical counterparts of Equations (1) and (2) may not be achieved after IPW, resulting in estimators with substantial bias and variance. To achieve balance when there is a lack overlap, some methods target a different estimand, $E[h(\bm{X})\tau(\bm{X})]$, which corresponds to a modified target population with improved overlap. For this modified estimand, Equations (1) and (2) must hold for $h(\bm{x})b(\bm{x})$. Li et al., (2018) propose one such method, overlap weights ($h(\bm{x})=e(\bm{x})(1-e(\bm{x})$), which are specifically designed to ensure that the weights exist even when positivity does not hold. By modifying the target population, overlap weights can more easily satisfy Equation (1) yielding an estimator with reduced bias and variance with respect to the modified estimand, the ATO. However, the overlap weights estimator tends to be biased for the ATE if there is treatment effect heterogeneity and the modified population is often challenging to characterize with respect to the original research question. Propensity score trimming (Crump et al.,, 2009) also modifies the target population to achieve improved overlap by discarding observations.
While the weights corresponding to these modified estimands tend to achieve improved finite-sample balance in scenarios with a lack of overlap, they rarely achieve exact finite-sample balance of empirical versions of Equations (1) and (2), which results in error in the corresponding treatment effect estimator. In fact, even when positivity does hold, exact finite-sample balance is rarely achieved by IPW-style estimators. While exact balance of covariate sample means is achieved for the ATO when the propensity score is estimated with a logistic regression (Li et al.,, 2018), for a majority of estimands, exact balance of the sample means will not be achieved, especially when there are positivity violations. This has motivated the development of sample weights that explicitly achieve three-way finite sample balance for a set of covariate functions.
Review of direct balancing weights
Direct balancing weights for the ATE achieve exact three-way balance for some pre-specified set of covariate functions; for the direct balancing weights estimator to be consistent for the ATE, $\mu_{z}(\bm{x})$ must be a linear combination of this set of covariate functions. Consider the case where we want to balance the covariate functions $b(\bm{x})=\{b_{j}(\bm{x})\}_{j=1}^{J}$. Then, direct balancing weights for the ATE take the following form,
$\{w^{bw}_{i}\}_{i=1}^{n}=\text{argmin}_{w_{i}}\sum_{i=1}^{n}D(w_{i})\>\>\>$ subject to (3a)
$\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]b_{j}(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}b_{j}(\bm{X}_{i})$ $=0\>\>\>j=1,\ldots,J,\>\>\>z=0,1$ (3b)
where $D(\cdot)$ is a convex measure of dispersion. The constraints $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ are often added to ensure that the weights do not extrapolate the observed outcomes. The resulting direct balancing weights estimator is
$\hat{\tau}_{w^{bw}}=\frac{1}{n}\sum_{i=1}^{n}w^{bw}_{i}Z_{i}Y_{i}-\frac{1}{n}\sum_{i=1}^{n}w^{bw}_{i}(1-Z_{i})Y_{i}.$ (4)
In this constrained optimization problem, constraint (3b) ensures that the treated and control covariate functions are balanced to the sample population. These constrains also imply that the treated and control covariate functions are balanced to each other (i.e., $\frac{1}{n}\sum_{i=1}^{n}w_{i}Z_{i}b_{j}(\bm{X}_{i})=\frac{1}{n}\sum_{i=1}^{n}w_{i}(1-Z_{i})b_{j}(\bm{X}_{i})$) such that three-way balance is satisfied by $\{w^{bw}_{i}\}_{i=1}^{n}$ for $b(\bm{x})$. Since the treated and control balance is implicitly satisfied, Problem (3) is separable by treatment group such that two separate constrained optimization problems can be solved to derive $\{w^{bw}_{i}\}_{i=1}^{n}$ (Chan et al.,, 2016); we discuss implications of this separation in Section 3.1.
When there is a lack of overlap between $\{b(\bm{x})\}_{z=1}$ and $\{b(\bm{x})\}_{z=0}$, often there is no solution to Problem (3) (i.e., Equations (3a)-(3b)) such that weights which achieve exact three-way balance cannot be derived. In these scenarios, the exact balance constraint is relaxed and approximate balance is achieved by constraining the absolute difference between sample means to be less than or equal to some parameter $\delta$; weights derived through this process are called minimal weights Wang and Zubizarreta, (2020). While the minimal weights estimator does have reduced variance compared to IPW and direct balancing estimators, minimal weights achieve none of the three balancing equations. Therefore, the minimal weights estimator may have bias due to both deviation from the target population and covariate imbalance. Thus, there is a need for methods that 1) achieve exact finite-sample covariate balance between treated and control distributions to reduce estimator error; and 2) modify the target population minimally and/or in an interpretable manner to yield an estimator with reduced variance when there is a lack of overlap.
A new partially retargeted balancing approach under positivity violations
To address these challenges, we propose a novel constrained optimization problem for deriving weights, primarily for scenarios where there is no solution to direct balancing weights due to positivity violations. Consider the case where Assumption Assumption 2: Positivity does not hold, but $0<\text{Pr}\{Z_{i}=1|c(\bm{X})=c(\bm{x})\}$ for some $c(\bm{x})\subset b(\bm{x})$ such that positivity holds when conditioning on this restricted subset. Let $g(\bm{x})$ be the remaining covariate functions in $b(\bm{x})$ such that $b(\bm{x})=[c(\bm{x})=\{c_{k}(\bm{x})\}_{k=1}^{K},g(\bm{x})=\{g_{l}(\bm{x})\}_{l=1}^{L}]$. The key idea of our method is to relax the constraints by balancing all covariate functions, $b(\bm{x})=\{c(\bm{x}),g(\bm{x})\}$, between treated and control groups, but only $c(\bm{x})$ to the sample population, yielding the following set of balancing equations:
$\{w^{*}_{i}\}_{i=1}^{n}=\text{argmin}_{w_{i}}\sum_{i=1}^{n}D(w_{i})\>\>\>$ subject to (5a)
$\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]c_{k}(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}c_{k}(\bm{X}_{i})$ $=0,\>\>\>k=1,\ldots K,\>\>\>z=0,1$ (5b)
$\sum_{i=1}^{n}w_{i}Z_{i}g_{l}(\bm{X}_{i})-\sum_{i=1}^{n}w_{i}(1-Z_{i})g_{l}(\bm{X}_{i})$ $=0,\>\>\>l=1,\ldots L.$ (5c)
The primary difference between this constrained optimization problem and the direct balancing weights constrained optimization problem is Equation (5c); in this equation, the functions, $g(\bm{x})=\{g_{l}(\bm{x})\}_{l=1}^{L}$ are only balanced between treated and control groups and not to the sample population. Thus, the derived weights satisfy exact three-way balance for functions $c(\bm{x})=\{c_{k}(\bm{x})\}_{k=1}^{K}$ but only balance between treated and control distributions for $g(\bm{x})$. In doing so, the derived weights do not balance $g(\bm{x})$ to a specific modified population (such as with overlap weights), however, the weighted treated group sample means of $g(\bm{x})$ are implicitly modified from the sample population mean when minimizing weight dispersion subject to only the treated and control balance constraint. We call these weights, $\{w_{i}^{*}\}_{i=1}^{n}$, partially retargeted balancing weights (PRTBW). In Section 3.1, we formally show that this approach is justified under a significantly relaxed positivity assumption. In addition, the corresponding estimator to Problem (5) (i.e., Equations (5a) - (5c)) is consistent for the ATE when $c(\bm{x})$ contains all treatment effect modifiers, which we discuss further in Section 3.2.
Figure 1: Simulated data exhibiting the different between direct balancing weights, minimal weights, and the proposed partially retargeted balancing weights. Squares (purple) and triangles (orange) indicate the unweighted control and treated sample means, respectively. The filled diamond indicates the sample mean while the outlined diamond indicates the weighted sample mean given $\{w_{i}^{*}\}$ derived through Problem (5). In the right-hand figures, the horizontal line indicates the sample mean while the vertical lines end at the weighted treated and control means for balancing weights (BW), minimal weights (MW) and our proposed weights (PRTBW). For this simulated data, there is no solution to Problem (3), exhibited by the dashed vertical line.
Figure 1 presents a toy example of the differences in covariate balance and target population between balancing weights (Problem (3)), minimal weights, and our proposed partially retargeted balancing weights (Problem (5)). For this simulated data, there is no solution to Problem (3), such that weights that exactly balance the treated and control sample means of $b(\bm{x})=\{c(\bm{x}),g(\bm{x})\}$ cannot be derived. Our proposed weights balance the treated and control sample means of $c(\bm{x})$ to the sample mean, but the treated and control sample means of $g(\bm{x})$ are balanced to a value that is substantially above the sample mean, providing a viable solution to Problem (5). In contrast, minimal weights do not achieve covariate balance for either $c(\bm{x})$ or $g(\bm{x})$. In fact, the weighted treated sample means are unchanged and for both $c(\bm{x})$ and $g(\bm{x})$ and the weighted control sample mean of $c(\bm{x})$ is further away from the sample mean of $c(\bm{x})$ than the unweighted control sample mean. Thus, while both the minimal weights optimization problem and our proposed Problem (5) yield solutions for this data, they do this through different mechanisms and each yields a solution with a different interpretation.
To further explore the theoretical properties of our proposed weights, we identify the dual formulation of Problem (5). The addition of constraint (5c) ensures that the optimization problem is not separable across the treatment groups such that there is correspondingly a single dual problem.
Theorem 1.
The dual of of Problem (5) is equivalent to the optimization problem
$\begin{split}\underset{\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}}{\textnormal{minimize}}\>\>\>&\frac{1}{n}\sum_{i=1}^{n}-Z_{i}\rho\{\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X}_{i})\}-(1-Z_{i})\rho\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X}_{i})\}\\ &+\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\alpha}_{0}^{T}c(\bm{X}_{i})\end{split}$ (6)
where $\bm{\alpha}_{0_{K\times 1}},\bm{\alpha}_{1_{K\times 1}},$ and $\bm{\gamma}_{L\times 1}$ are the dual variables associated with the balancing constraints in Equations (5b)-(5c), $\rho(t)=t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(1)$ is strictly concave, and h(t)=D(1t)h(t)=D(1-t). Furthermore, the primal solution $w_{i}^{*}=Z_{i}w^{*}_{1i}+(1-Z_{i})w^{*}_{0i}$ satisfies $w^{*}_{1i}=\rho^{\prime}\{\hat{\bm{\alpha}}_{1}^{T}c(\bm{X}_{i})+\hat{\bm{\gamma}}^{T}g(\bm{X}_{i})\}\>\>\text{and}\>\>w^{*}_{0i}=\rho^{\prime}\{\hat{\bm{\alpha}}_{0}^{T}c(\bm{X}_{i})-\hat{\bm{\gamma}}^{T}g(\bm{X}_{i})\}$ where $\hat{\bm{\alpha}}_{0},\hat{\bm{\alpha}}_{1},\hat{\bm{\gamma}}$ is the solution to the dual problem.
The proof of this theorem is in supplementary Section 1. In the following sections, we discuss the implications of Problem (5) in terms of both positivity and the interpretation of the resulting estimator while drawing comparisons to direct balancing weights, minimal weights, and overlap weights.
Positivity considerations
The construction of Problem (5) is primarily motivated by the challenges posed by lack of overlap on causal effect estimation, generally, and direct balancing weights, specifically. In the following section, we show that Problem (5) corresponds to a relaxed positivity assumption where there is a solution to Problem (5) even when the standard positivity assumption does not hold. Furthermore, we discuss the restrictions placed on the implied propensity score model of Problem (5) and how this restriction connects to the relaxed positivity assumption. We additionally connect these results to the positivity assumptions and implied propensity score models of direct balancing, minimal, and overlap weights.
While positivity is required for the existence of a solution to Problem (3) (Zhao and Percival,, 2017), by not balancing $g(\bm{x})$ to the target population, the weights derived through Problem (5) exist under a relaxed positivity assumption:
Assumption 2: Relaxed positivity.
$0<\text{Pr}\{Z_{i}|c(\bm{X}_{i})=c(\bm{x})\}<1$ for all $\bm{x}$ and there exists an $\bm{x}^{*}$ such that $0<\text{Pr}\{Z_{i}|c(\bm{X}_{i})=c(\bm{x}^{*}),g(\bm{X}_{i})=g(\bm{x}^{*})\}<1$.
The following proposition states that this assumption is sufficient for the existence of the weights in Problem (5) as $n\to\infty$.
Proposition 2.
Suppose Assumption Assumption 2: Relaxed positivity is satisfied and the expectation of $c(\bm{X})$ exist, then $P(w^{*}\textnormal{ exists})\to 1$ as $n\to\infty$.
We leave the proof of this proposition to supplementary Section 1, however here we provide a brief intuitive discussion of how Problem (5) often yields a solution in finite samples when Problem (3) does not. The weights in Problem (3) exist if the convex hulls generated by $\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=1}$ and $\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=0}$ contain $\frac{1}{n}\sum_{i=1}^{n}\{c(\bm{X}_{i}),g(\bm{X}_{i})\}$ (Zhao and Percival,, 2017). In contrast, the weights in Problem (5) exist if 1) the convex hulls of $\{c(\bm{X}_{i})\}_{Z_{i}=1}$ and $\{c(\bm{X}_{i})\}_{Z_{i}=0}$ contain $\frac{1}{n}\sum_{i=1}^{n}c(\bm{X}_{i})$; and 2) $\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=1}\cap\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=0}\neq\emptyset$. Thus, as long as any $J$ dimensional vector is in the intersection of the convex hulls, rather than the $J$ dimensional sample mean, the existence of a solution is guaranteed. This is a substantial relaxation of the requirements for weight existence, especially for scenarios where one treatment group is much smaller than the other. Returning to Figure 1, the sample mean of $\{c(\bm{x}),g(\bm{x})\}$ is not in the intersection of the two convex hulls and thus there is no solution for direct balancing weights as previously discussed. However, there are many observations in the intersection of the two convex hulls. Further, when examining $c(\bm{x})$, its sample mean is within the observed range of $c(\bm{x})$ values for both treatment groups. Thus, there is a solution to our proposed Problem (5), and in fact the corresponding weighted sample mean given by $\{w_{i}^{*}\}_{i=1}^{n}$ is within the intersection of the treated and control convex hulls.
The relaxed positivity assumption for Problem 5 is directly related to the propensity score model implied by Problem 5. Taking the expectation of population version of the loss function in Equation (6) with respect to $\bm{x}$ yields the following first order conditions,
$\{e(\bm{x})\}^{-1}$ $=\rho^{\prime}\{\bm{\alpha}_{1}^{T}c(\bm{x})+\bm{\gamma}^{T}g(\bm{x})\}=w_{1}^{*},$ (7a)
$\{1-e(\bm{x})\}^{-1}$ $=\rho^{\prime}\{\bm{\alpha}_{0}^{T}c(\bm{x})-\bm{\gamma}^{T}g(\bm{x})\}=w_{0}^{*},\>\>\>\text{and}$ (7b)
$e(\bm{x})$ $=\frac{\rho^{\prime}\{\bm{\alpha}_{0}^{T}c(\bm{x})-\bm{\gamma}^{T}g(\bm{x})\}}{\rho^{\prime}\{\bm{\alpha}_{0}^{T}c(\bm{x})-\bm{\gamma}^{T}g(\bm{x})\}+\rho^{\prime}\{\bm{\alpha}_{1}^{T}c(\bm{x})+\bm{\gamma}^{T}g(\bm{x})\}}=\frac{w_{0}^{*}}{w_{0}^{*}+w_{1}^{*}}.$ (7c)
Thus, the weights $w_{0}^{*}$ and $w_{1}^{*}$ correspond to IPW for a specific propensity score model, which we reference as the implied propensity score model of Problem (5). In this implied propensity score model, the dual variables $\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}$ can be considered as the model coefficients. Furthermore, $\rho^{\prime}(\cdot)$ may be regarded as the link function of the generalized linear propensity score model, where the selected measure of dispersion (i.e., $D(\cdot)$) determines this link function (Wang and Zubizarreta,, 2020).
Problem (5) restricts the coefficients corresponding to $g(\bm{x})$ in the models for $\{e(\bm{x})\}^{-1}$ and $\{1-e(\bm{x})\}^{-1}$ to be opposites of each other. This ensures that $w^{*}_{0}=\{1-w_{1}^{*^{-1}}\}^{-1}$ such that only a single propensity score model is specified by Problem (5). However, there there are substantial restrictions on the propensity score model that satisfies conditions (7a)-(7c). Specifically, $\text{Var}(Z|\bm{X}=\bm{x})$ is restricted to be a function of only $c(\bm{x})$ for specific dispersion measures (see supplementary Section 2 for further discussion). Since this generally does not hold when $E[Z|\bm{X}=\bm{x}]$ is a function of both $c(\bm{x})$ and $g(\bm{x})$, this model will only hold when $g(\bm{x})$ takes a specific functional form. However, it is explicitly through this restriction on the propensity score model that there is a solution to Problem (5) in scenarios with a lack of overlap, as overlap is improved through a reduction in the ability to predict treatment assignment (Clivio et al.,, 2024). Rather than attempt to specify a correct implied propensity score model, Problem (5) imposes restrictions on the propensity score model to improve overlap and yield a solution to the optimization problem.
Connections to other methods
Problem (5) corresponds to a single propensity score model, while Problem (3) implies two incompatible propensity score models due to the separability of the optimization problem by treatment group. By implying a single propensity score model, rather than two incompatible propensity score models, the weights derived by our proposed method maintain the standard philosophical interpretation of inverse probability weights that direct balancing weights lack for the ATE. Furthermore, this allows for the doubly robust property (i.e., the estimator is consistent if either the implied propensity score model or outcome model is properly specified) and semiparametric efficiency of our method given the correct model specification that direct balancing weights for the ATE do not satisfy (see Section 4 for further discussion of these properties). Minimal weights also correspond to two incompatible propensity score models but place an l1 penalty on the dual variables (i.e., propensity score model coefficients). While the restrictions imposed on the implied propensity score model of Problem (5) are comparably less intuitive, by restricting $\text{Var}(Z|\bm{X}=\bm{x})$ rather than $E[Z|\bm{X}=\bm{x}]$ our derived weights are able to achieve exact finite sample balance, unlike minimal weights. Furthermore, reducing the variance of the derived weights is key to reducing estimator variance, which restricting $\text{Var}(Z|\bm{X}=\bm{x})$ yields; thus, our proposed method precisely restricts the aspect the propensity score model that primarily influences estimator variance.
When $c(\bm{x})=\emptyset$ and $g(\bm{x})=b(\bm{x})$ in Problem (5), Assumption Assumption 2: Relaxed positivity becomes: “There exists an $\bm{x}^{*}$ such that $0<\text{Pr}\{Z_{i}|c(\bm{X}_{i})=c(\bm{x}^{*}),g(\bm{X}_{i})=g(\bm{x}^{*})\}<1$” which is comparable to the lack of a positivity assumption for overlap weights. Thus, Assumption Assumption 2: Relaxed positivity is an intermediate between Assumption $2$ and no positivity assumption; as more covariate functions are added to the $g(\bm{x})$ set, Assumption Assumption 2: Relaxed positivity becomes more relaxed. To further connect overlap weights and Problem (5), when $c(\bm{x})=\emptyset$ and $g(\bm{x})=b(\bm{x})$ the first order conditions in (7) simplify to $e(\bm{x})/\{1-e(\bm{x})\}=w_{0}^{*}/w_{1}^{*}$. While the exact solution to the dual problem in this case will depend on the dispersion measure, overlap weights satisfy this condition. Therefore, Problem (5) with $c(\bm{x})=\emptyset$ and $g(\bm{x})=b(\bm{x})$ can be considered the direct balancing equivalent of overlap weights; however, Problem (5) will always yield exact balance while overlap weights only yields exact balance when the propensity score is estimated with a logistic regression.
Interpretation of the proposed estimator
While Problem (5) yields a solution when there is a lack of overlap, in doing so the population of $g(\bm{x})$ is modified away from the ATE target population, potentially resulting in estimator bias with respect to the ATE (termed bias due to estimand mismatch in Barnard et al., (2025)). In the following sections, we provide model- and design- based perspectives on the conditions under which our proposed estimator has minimal error with respect to the ATE. In addition, we provide a characterization of the modified population and estimand that $\tau_{w^{*}}$ targets when $\tau_{w^{*}}$ does not have minimal error with respect to the ATE. To discuss when our proposed estimator has minimal error with respect to the ATE, we first explore the following error decomposition for a normalized weighted estimator for the ATE,
$\hskip-14.45377pt\hat{\tau}_{w}-\tau$ $=\frac{1}{n}\sum_{i=1}^{n}\{w(\bm{X}_{i})Z_{i}\mu_{Z_{i}}(\bm{X}_{i})-\mu_{1}(\bm{X}_{i})\}-\frac{1}{n}\sum_{i=1}^{n}\{w(\bm{X}_{i})(1-Z_{i})\mu_{Z_{i}}(\bm{X}_{i})-\mu_{0}(\bm{X}_{i})\}$ (8a)
$-\int\{\mu_{1}(\bm{x})-\mu_{0}(\bm{x})\}d\{F-F_{n}\}(\bm{x})$ (8b)
$+\frac{1}{n_{1}}\sum_{i=1}^{n}w_{1}(\bm{X}_{i})\epsilon_{i}Z_{i}-\frac{1}{n_{0}}\sum_{i=1}^{n}w_{0}(\bm{X}_{i})\epsilon_{i}(1-Z_{i}),$ (8c)
where $\epsilon_{i}=Y_{i}(Z_{i})-\mu_{Z_{i}}(\bm{X}_{i})$ and $F_{n}(\bm{x})=\sum_{i=1}^{n}I(\bm{X}_{i}\leq\bm{x})/n$ is the empirical CDF. In this expression, term (8b) goes to zero for a representative sample of the population and term (8c) has expectation zero. Thus, term (8a) is the primary contributor to estimator error; we focus on this term for the remainder of this section.
To better understand the conditions under which our balancing equations are justified, consider the case where $\mu_{z}(\bm{x})$ is a linear function of $c(\bm{x})$ and $g(\bm{x})$ such that $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}_{1}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}_{0}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}_{z}\in\mathbb{R}^{L}$. When weights are derived through Problem (3), term (8a) will be zero for the corresponding estimator. However when weights are derived with our proposed Problem (5), term (8a) simplifies to $\frac{1}{n}\sum_{i=1}^{n}\bm{\lambda}_{1}^{T}g(\bm{X}_{i})-\bm{\lambda}_{0}^{T}g(\bm{X}_{i})$ implying that term (8a) is zero when $\bm{\lambda}_{1}=\bm{\lambda}_{0}$. Thus, term (8a) for our proposed estimator is zero when $\mu_{z}(\bm{x})=\bm{\beta}_{z}^{T}c(\bm{x})+\bm{\lambda}^{T}g(\bm{x})$ and $\tau(\bm{x})=(\bm{\beta}_{1}^{T}-\bm{\beta}_{0}^{T})c(\bm{x})$. These models imply that our proposed estimator has minimal error with respect to the ATE when the set $c(\bm{x})$ contains all treatment effect modifiers. Thus, as the set $c(\bm{x})$ gets smaller and correspondingly the set $g(\bm{x})$ gets larger, our proposed estimator requires a stronger assumption about the number of effect modifiers to be justified for the ATE. However, this is a relaxation of the assumption $\tau(\bm{x})=\tau$ required for the overlap weights estimator to be unbiased for the ATE.
We can alternatively explore term (8a) for our proposed estimator from a design-based, rather than modeling, prospective; an alternative simplification of term (8a) is $(\bm{\lambda}_{1}^{T}-\bm{\lambda}_{0}^{T})\left\{\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]g(\bm{X}_{i})\right$. $\left.-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})\right\}$. From a design-based perspective, this term tends to be small when $1^{T}\left\lvert\frac{1}{n}\sum_{i=1}^{n}w^{*}(\bm{X}_{i})I[Z_{i}=z]g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})\right\rvert$ is small. However, in Problem (5) we explicitly do not constrain $\frac{1}{n}\sum_{i=1}^{n}w^{*}(\bm{X}_{i})I[Z_{i}=z]g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})=0$ in order to yield a solution when there is a lack of overlap. Thus, $1^{T}\left\lvert\frac{1}{n}\sum_{i=1}^{n}w^{*}(\bm{X}_{i})I[Z_{i}=z]g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})\right\rvert$ will be non-zero when weights are derived using Problem (5). Despite this, the error of our proposed estimator with respect to the ATE is due to differences between the weighted and sample populations of $g(\bm{x})$, rather than the imbalance between the weighted treated and control populations. Thus, when $c(\bm{x})$ does not contain all treatment effect modifiers, the estimator has minimal error with respect to the average treatment effect estimand over a modified population of $g(\bm{x})$, which can be characterized by $\frac{1}{n}\sum_{i=1}^{n}w_{i}Z_{i}g(\bm{X}_{i})=\frac{1}{n}\sum_{i=1}^{n}w_{i}(1-Z_{i})g(\bm{X}_{i})$ in comparison to the overall sample mean. While the ATO population can also be characterized through weighted sample means, it can be challenging to conceptualize this modified population especially with high-dimensional covariates. In contrast to both overlap weights and our proposed estimator, the minimal weights estimator does not target another estimand of interest and its bias with respect to the ATE is predominantly due to covariate imbalance. We provide further guidance and discussion on how to select $g(\bm{x})$ given these model- and design-based perspectives in Section 6 and supplementary Section 4.1.
Asymptotic properties of the proposed estimator
We derive asymptotic results of the solution to Equation (6) which we then use to obtain asymptotic results of our proposed estimator, $\tau_{w^{*}}$, under various conditions. All proofs of results in this section are in supplementary Section 1. We first require an additional regularity condition, similar to those in Hirano et al., (2003); Chan et al., (2016) and Källberg and Waernbaum, (2023), on the set balancing functions, $b(\bm{x})$.
Assumption 3: Balance function regularity conditions.
There exists a finite $M$, such that $|b_{j}(\bm{X})|<M$ for $j=1,\ldots,J$ almost surely and $B(\bm{X})=[b_{1}(\bm{X}),\ldots,b_{J}(\bm{X})]$ has full rank.
For the remainder of this section, we notate $w_{i}^{*}$ as $w^{*}(\bm{X}_{i})$ to explicitly show the relationship to $\bm{X}_{i}$. Given Assumptions Assumption 1: Strong ignorability, Assumption 2: Relaxed positivity, and Assumption 3: Balance function regularity conditions, we show that $w_{1}^{*}(\bm{X}_{i})$ and $w_{0}^{*}(\bm{X}_{i})$ are consistent for their asymptotic counters, $\tilde{w}_{1}(\bm{X}_{i})$ and $\tilde{w}_{0}(\bm{X}_{i})$ that uniquely solve the population version of Equation 6 (supplementary Section 1, Proposition 5). Since Problem (5) implies a single propensity score model, our proposed estimator is doubly robust as stated below in Theorem 3; in contrast, $\hat{\tau}_{w^{bw}}$ is not doubly robust (Chan et al.,, 2016).
Theorem 3.
Suppose Assumptions Assumption 1: Strong ignorability and Assumption 3: Balance function regularity conditions hold. Assume that the expectation of $c(\bm{x})$ exists and that $\text{Var}\{Y(a)\}<\infty$ for $a=0,1$. Then our proposed estimator, $\hat{\tau}_{w^{*}}$, is doubly robust in the sense that if either 1) Assumption Assumption 2: Positivity holds and $e(\bm{x})$ satisfies conditions (7a)-(7c) or; 2) Assumption Assumption 2: Relaxed positivity holds, $\mu_{z}(\bm{x})$ is linear in $c(\bm{x})$ and $g(\bm{x})$, and $\tau(\bm{x})$ is linear in $c(\bm{x})$ then $\hat{\tau}_{w^{*}}$ is consistent for the ATE.
While $\hat{\tau}_{w}^{*}$ is doubly robust, we note that $e(\bm{x})$ will rarely satisfy the restrictive conditions (7a)-(7c). Still, our proposed estimator is consistent under the relaxed positivity assumption if $c(\bm{x})$ contains all effect modifiers. When $c(\bm{x})$ does not contain all effect modifiers, our proposed estimator is consistent for the modified estimand $\tau_{g,\tilde{w}}=E[E[Y|Z=1,c(\bm{X}),\tilde{w}_{1}(\bm{X})e(\bm{X})g(\bm{X})]]-E[E[Y|Z=0,c(\bm{X}),\tilde{w}_{0}(\bm{X})\{1-e(\bm{X})\}g(\bm{X})]]$ as stated in Corollary 3.1.
Corollary 3.1.
Suppose Assumptions Assumption 1: Strong ignorability and Assumption 2: Relaxed positivity hold. Assume that the expectation of $c(\bm{x})$ exists and that $\text{Var}\{Y(a)\}<\infty$ for $a=0,1$. Then our proposed estimator, $\hat{\tau}_{w^{*}}$ is consistent for $\tau_{g,\tilde{w}}$ if $e(\bm{x})$ does not satisfy conditions (7a)-(7c) and $\mu_{z}(\bm{x})$, $\tau(\bm{x})$ are linear in $c(\bm{x})$ and $g(\bm{x})$.
Note that $\tau_{g,\tilde{w}}$ is indeed a causal estimand because $E[\tilde{w}_{1}(\bm{X}_{i})e(\bm{X}_{i})g(\bm{X_{i}})]=E[\tilde{w}(\bm{X}_{i})_{0}\{1-e(\bm{X}_{i})\}g(\bm{X_{i}})]$ by $\tilde{w}$ satisfying the first order conditions of the population version of Equation (6). Thus, we can also write $\tau_{g,\tilde{w}}=E[E[Y|Z=1,c(\bm{X}),\tilde{w}_{1}(\bm{X})e(\bm{X})g(\bm{X})]]-E[E[Y_{i}|Z_{i}=0,c(\bm{X}),\tilde{w}_{1}(\bm{X})e(\bm{X})g(\bm{X})]]$.
When positivity does hold and the implied propensity score and outcome models of Problem 5 are correctly specified, our proposed estimator is semiparametrically efficient as stated in Theorem 4.
Theorem 4.
Suppose Assumptions 11, 22, and 33 hold and $\text{Var}\{Y(a)|\bm{X}=\bm{x}\}<\infty$ for all $\bm{x}$ and $a=0,1$. Suppose $e(\bm{x})$ satisfies conditions (7a)-(7c), $\mu_{z}(\bm{x})$ is linear in $c(\bm{x})$ and $g(\bm{x})$, and $\tau(\bm{x})$ is linear in $c(\bm{x})$. Then
$\sqrt{n}(\hat{\tau}_{w^{*}}-\tau)\to N(0,V_{opt})$
where $V_{opt}=E\left[\frac{\text{Var}(Y(1)|\bm{X})}{e(\bm{X})}+\frac{\text{Var}(Y(0)|\bm{X})}{1-e(\bm{X})}+\{\tau(\bm{X})-\tau\}^{2}\right]$ is the semiparametric efficiency bound (Hahn,, 1998).
However, it is rare that the implied propensity score model is correctly specified and our proposed estimator is motivated by scenarios where Assumption Assumption 2: Positivity does not hold. We thus derive and present the asymptotic variance of our estimator when only the outcome models are correctly specified in Corollary 4.1.
Corollary 4.1.
Suppose Assumptions 11, $2^{*}$, and 33 hold and $\text{Var}\{Y(a)|\bm{X}=\bm{x}\}<\infty$ for all $\bm{x}$ and $a=0,1$. Suppose $e(\bm{x})$ does not satisfy conditions (7a)-(7c), $\mu_{z}(\bm{x})$ is linear in $c(\bm{x})$ and $g(\bm{x})$, and $\tau(\bm{x})$ is linear in $c(\bm{x})$. Then,
$\sqrt{n}(\hat{\tau}_{w^{*}}-\tau)\to N(0,V_{opt}^{*})$
where $V_{opt}^{*}=E[\tilde{w}_{1}(\bm{X})^{2}e(\bm{X})\text{Var}\{Y(1)|\bm{X}_{i}\}+\tilde{w}_{0}(\bm{X})^{2}\{1-e(\bm{X})\}\text{Var}\{Y(0)|\bm{X}_{i}\}+\{\tau(\bm{X})-\tau\}^{2}]$
Of key importance, this variance result only requires the relaxed positivity assumption, rather than the standard positivity assumption. However, to compare $V_{opt}$ and $V^{*}_{opt}$, consider the case where positivity does hold. Since the asymptotic version of the weights tend to be less extreme (i.e., $\tilde{w}_{1}(\bm{X}_{i})\geq e(\bm{X}_{i})$ and $\tilde{w}_{0}(\bm{X}_{i})\leq 1-e(\bm{X}_{i})$) than IPW through the restriction on the implied propensity score models, generally $V^{*}_{opt}\leq V_{opt}$. Thus, our proposed estimator does indeed have reduced asymptotic variance, especially in scenarios where $e(\bm{x})$ is close to zero or positivity does not hold. While intuitively as the $g(\bm{x})$ set gets larger, estimator variance decreases, Corollary 4.1 formalizes this intuition; as the $g(\bm{x})$ set gets larger, $\tilde{w}_{z}(\bm{x})$ become less extreme and correspondingly $V^{*}_{opt}$ becomes smaller.
Related estimation problems
Our proposed constrained optimization problem for deriving weights for an ATE estimator can be readily extended to variety of additional estimation problems. We discuss extensions to the ATT and transporting effects below. Discussion of causal effect estimation with more than two treatment groups, weighted average treatment effects, and distributional balancing weights are in supplementary Section 3.
ATT
When deriving either IPW or direct balancing weights for the ATT, only weights for the control group are derived such that the corresponding estimator is $\frac{1}{n}\sum_{i=1}^{n}Y_{i}-\frac{1}{n}\sum_{i=1}^{n}w_{i}(1-Z_{i})Y_{i}$. However, when adapting Problem (5) for ATT estimation, weights must be derived for both the treated and control group. To derive these weights, $\{w_{i}^{*ATT}\}$, we can simply replace the right-hand side of constraint (5b) with $\frac{1}{n}\sum_{i=1}^{n}Z_{i}c_{k}(\bm{X}_{i})$ and compute Equation (4) with these weights. The design-based and model-based interpretations of the sets $c(\bm{x})$ and $g(\bm{x})$ remain the same as described in Section 3.2. However, the covariates that may be effect modifiers for the entire sample population potentially could not be for the treated population. Further, the methods for a data-driven selection of $g(\bm{x})$ proposed in Section 6 can be used here as well.
Transporting effects
While clinical trials yield ‘gold standard’ causal effect estimates through randomization, the inclusion/exclusion trial criteria is often restrictive such that the trial population may not reflect a population of interest. In these cases, it is desirable to transport the effect to a more meaningful population. However, there is often a lack of overlap between the restrictive trial population and the target population such that methods similar to propensity score trimming or overlap weights must be used. Our procedure allows the analyst to improve overlap by modifying the target population for only a subset of covariates, rather than modifying all covariates or discarding observations, which is inconsistent with the goal of estimating the effect for a meaningful population. Before describing the adaptation of Problem 5 for estimating a transported effect, we introduce some additional notation. Let $R_{i}=r$ be an indicator of whether an observation belongs to the trial population ($r=1$) or the target population ($r=0$). We only observe $Y_{i}$ and $Z_{i}$ when $R_{i}=1$ but observe $\bm{X}_{i}$ for both populations. Let $n_{r}$ indicate the sample size in population $r$. Given additional assumptions outlined in Dahabreh et al., (2020), the transported effect estimand is identifiable with observed data. We propose the following modification of Problem (5) for transportability,
$\{w^{*t}_{i}\}_{i=1}^{n}=$ $\text{argmin}_{w_{i}}\sum_{i=1}^{n}D(w_{i})\>\>\>\text{subject to}$ (9a)
$\frac{1}{n_{1}}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]R_{i}c_{k}(\bm{X}_{i})$ $=\frac{1}{n_{0}}\sum_{i=1}^{n}(1-R_{i})c_{k}(\bm{X}_{i}),\>\>\>k=1,\ldots K,\>\>\>z=0,1$ (9b)
$\frac{1}{n_{1}}\sum_{i=1}^{n}w_{i}Z_{i}R_{i}g_{l}(\bm{X}_{i})$ $=\frac{1}{n_{1}}\sum_{i=1}^{n}w_{i}(1-Z_{i})R_{i}g_{l}(\bm{X}_{i}),\>\>\>l=1,\ldots L$ (9c)
Then, our proposed estimator for the transported effect is $\tau_{w^{*t}}=\frac{1}{n}\sum_{i=1}^{n}w^{*t}_{i}Z_{i}R_{i}Y_{i}-\frac{1}{n}\sum_{i=1}^{n}w^{*t}_{i}(1-Z_{i})R_{i}Y_{i}$. The interpretation of $c(\bm{x})$ and $g(\bm{x})$ remains the same as described in Section 3.2 however the selection of $g(\bm{x})$ may be of even more importance as maintaining as much of the target population as possible is of key interest when transporting effects.
Implementation details
To select the optimal $g(\bm{x})$ set, in terms of error with respect to the ATE, one approach is to examine all possible combinations of functions of covariates in the set $g(\bm{x})$ and select the set that 1) yields a solution to Problem (5); and 2) minimizes term (8a). However, examining all possible covariate subsets is computationally prohibitive, especially as covariate dimension increases. Therefore, we propose a greedy algorithm (Algorithm 1) for identifying an approximately optimal $g(\bm{x})$ subset according to a specific metric of interest aligned with either the model- or design- based perspective on estimator error. We also propose an additional algorithm (supplementary Algorithm 1) that is less computationally intensive than Algorithm 1; we found these algorithms resulted in estimators with similar performance when there was no correlation between covariates (supplementary Figure 2). However, we generally recommend the use of Algorithm 1 if it is not computationally prohibitive.
Algorithm 1 Identifying an approximately optimal $g(\bm{X})$ set given metric $m$
1:Inputs: Set of functions of covariates $b_{1}(\bm{X}),\ldots,b_{J}(\bm{X})$ and a metric $m$ that can be computed for each $b_{j}(\bm{X})$
2:Initialize: $b(\bm{X})^{-}=\emptyset$
3:while $w^{*}$ given $g(\bm{X})=b(\bm{X})^{-}$ does not exist do
4:  Compute metric $m$ for $\{b_{1}(\bm{X}),\ldots,b_{j}(\bm{X})\}\setminus b(\bm{X})^{-}\}$ and identify $b^{*}(\bm{X})$, the covariate with the minimum or maximum value of the metric $m$
5:  Let $b(\bm{X})^{-}=\{b(\bm{X})^{-},b^{*}(\bm{X})\}$
6:  Solve for $w^{*}$ in Problem (5) for $g(\bm{X})=b(\bm{X})^{-}$ and $c(\bm{X})=\{b_{1}(\bm{X}),\ldots,b_{j}\}\setminus b(\bm{X})^{-}$
7:end while
8:Outputs:
9:   $b(\bm{X})^{-}$ $\triangleright$ Smallest set (given metric $m$) that yields a solution to Problem (5) when $g(\bm{X})=b(\bm{X})^{-}$
10:   $\{w^{*}\}_{i=1}^{n}$ $\triangleright$ Resulting weights from (5) with $g(\bm{X})=b(\bm{X})^{-}$
From a design-based perspective, selecting the covariate functions that are each conditionally the most strongly predictive of treatment will tend to result in the smallest $g(\bm{x})$ set that yields a solution to Problem (5). While selecting this set may not minimize the design-based components of estimator error, this set does tend to result in smaller design-based estimator error and a simpler interpretation of the resulting estimator. Thus, for the design-based metric, we propose and implement the Spearman semipartial correlation; in step four the maximum correlation is used to identify $b^{*}(\bm{x})$. We choose this metric because it is rank-based and reflects the impact of removing a covariate in a hypothetical propensity score model, however other metrics could work well here. Our design-based estimator takes the same form as Equation (4) with weights computed with the $g(\bm{x})$ set selected by Algorithm 1.
From a model-based perspective, selecting the covariate functions that do not contribute to treatment effect heterogeneity as the $g(\bm{x})$ set will tend to minimize estimator error. Thus, we implement the absolute value of the one-step doubly robust estimator of treatment effect modification, proposed by Boileau et al., (2025), as the model-based metric to use within Algorithm 1; in step four the minimum of this metric is used to identify $b^{*}(\bm{X})$. Our model-based estimator is a cross-fit style estimator to preserve downstream inference. We provide further discussion about this metric and model-based estimator in supplementary Section 4.2. For either metric/estimator, Algorithm 1 can be initiated with a non-empty $b(\bm{x})^{-}$ set (step 2) if there are covariates that one wants to include in the $g(\bm{x})$ set a priori. If minimizing error with respect to the ATE is not a priority, $g(\bm{x})$ could be chosen such that the modified $g(\bm{x})$ population is of scientific interest which we discuss further in supplementary Section 4.1.
Simulation experiments
We generate a variety of simulation scenarios that vary in the number of confounders (i.e., covariates), level of overlap, level of treatment effect heterogeneity, and proportion of covariates that are effect modifiers. Through these scenarios, we first aim to validate the estimator corresponding to the weights derived through Problem (5) when the true $\tau(\bm{x})$ function is known. In addition, we seek to determine in which data generating scenarios the proposed methods for a completely data-driven selection of $g(\bm{x})$ in Section 6 yield estimators with improved performance when compared to IPW ATE estimation, minimal weights, and the ATO when there is no solution to Problem (3).
Data generation
For clarity in presentation, we drop the observation index on the variables. For $p=20$ and $p=100$ we simulate $V_{1}$ - $V_{p}$ from a truncated normal distribution where $E[V_{i}]=0$, $\text{Var}(V_{i})=1$ and $\text{Cov}(V_{i},V_{j})=0$ for all $i=1,\ldots,p$ and $i\neq j$. To obtain a mix of continuous and categorial coefficients, we take $X_{1}$ - $X_{p/2}=$ $V_{1}$ - $V_{p/2}$ and $X_{j}=I[V_{j}<0]$ for $j=p/2+1,\ldots p$. We generate the true propensity scores with a logistic model, $e(\bm{X})=\{1-\exp(-\sum_{i=0}^{p}\alpha_{i})\}^{-1}$. In this model, $\alpha_{0}$ is selected to obtain the desired percent of treated observations (20%, 40%) and all other coefficients are determined as described in supplementary Section 5. We simulate treatment assignment as $Z\sim\text{Bernoulli}\{e(\bm{X})\}$ and the outcome as $Y=\mu_{z}(\bm{X})+N(0,1)$ for $Z=z$. Both $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ are modeled as linear combinations of $X_{1},\ldots,X_{p}$. For $\theta=25\%,75\%$ of covariates, their corresponding $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ coefficients are the equal such that these covariates are not effect modifiers. For all other covariates, the $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ coefficients are determined as described in supplementary Section 5
We generate a total of 72 simulation scenarios that are all combinations $p=20,100$, 20% and 40% treated, 25% and 75% of covariates being effect modifiers, three levels of overlap, and three levels of treatment effect heterogeneity. For each simulation scenario we generate 1,000 independent datasets of size either 1,000 ($p=20$) or 2,000 ($p=100$) observations. We calculate the true value of the ATE using 10 million Monte Carlo samples. We compute 1) IPW, direct balancing weights, and minimal weights estimators for the ATE; 2) our proposed estimator given the true $g(\bm{X})$ and the design- and model-based estimators proposed in Section 6; and 3) the IPW and balancing weights (i.e., all covariates in the set $g(\bm{X})$) estimators for the ATO. All balancing weights are computed with the stable balancing weights deviance measure (Zubizarreta,, 2015). We compare the mean squared error (MSE) of these estimators with respect to the true ATE.
Results
Our proposed estimator implemented with the true $\tau(\bm{x})$ function has the minimum estimator MSE with respect to the ATE across almost all simulation scenarios where there is a solution to Problem (5) for the true $\tau(\bm{x})$ function (Figure 2 and supplementary Figure 4). Specifically, this estimator tends to be minimally biased and have comparable variance to the minimal weights estimator and ATO estimator (supplementary Figures 5-8). Thus, when there is a lack of overlap and no solution to Problem (3), our proposed estimator is the only estimator that has minimal bias for the ATE; both the minimal weights and the ATO estimators tend to be substantially biased for the ATE. However, it is rare that the true $\tau(\bm{x})$ is known, so for the remainder of this section we focus on performance of our proposed design- and model-based estimators.
Figure 2: Estimator mean squared error (MSE) for all 20% treated simulation scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while each panel is labeled by different overlap levels, number of predictors, and percent of covariates that are effect modifiers.
Across almost all simulation scenarios, either our proposed design- or model-based estimator has minimum estimator MSE with respect to the ATE compared to minimal weights or either ATO estimator (Figure 2 and supplementary Figure 4). Furthermore, the design-based estimator has lower MSE than the IPW ATE estimator across all scenarios, which none of the other estimators achieve. The minimal weights estimator tends to have the largest MSE, even larger than ATE and ATO estimators, when there is moderate-high treatment effect heterogeneity and low overlap. The model-based estimator tends to perform better than the design-based estimator for $p=20$, in terms of MSE, while the reverse is true when $p=100$. We likely see this trend because there may not be sufficient sample size for estimating treatment effect modification and computing the weights with sample splitting. Thus, the model-based estimator requires a substantially larger sample size for high dimensional covariates for comparable performance to the design-based estimator. We see similar trends in estimator MSE between 20% treated (Figure 2) and 40% treated scenarios (supplementary Figure 4), though for 40% treated scenarios all estimators substantially outperform the IPW estimator in terms of MSE.
The scenarios where the minimal weights estimator tends to have lower MSE than either of our proposed estimators are ones with lower treatment effect heterogeneity (in terms of both the scale of covariate-level treatment effect heterogeneity and the % of covariates that are effect modifiers). In these scenarios, the design- and model-based estimators tend to have larger estimator variance, but smaller statistical bias with respect to the ATE compared to the minimal weights estimator (supplementary Figure 5). While the small variance of the minimal weights estimator is desirable, the estimator is substantially biased for the ATE such that the corresponding 95% CI may not include the true ATE. Furthermore, there is no clear causal estimand for which the minimal weights estimator is consistent. Thus, even in scenarios in which the minimal weights estimator has lower MSE, our proposed estimators may be preferable.
EHR study on the effect of indwelling arterial catheters on mortality
We apply our weighting method for ATE estimation to observational data from the MIMIC-III v1.4 critical care database (Johnson et al.,, 2016). Specifically, we perform a reanalysis of Hsu et al., (2015) which examines effect of indwelling arterial catheters (IACs) on mortality in patients with respiratory failure. The data can be obtained with queries provided at the GitHub repository https://github.com/MIT-LCP/mimic-code. The data has 2,522 observations of mechanically ventilated patients, where 51.5% of patients received IAC. The outcome of interest is an indicator of mortality 28 days within hospital admission. Pre-treatment covariates that may be confounders include demographic information, baseline measurements (e.g., blood pressure, lab values), risk scores, and more. Missing data is imputed with single imputation and we include missing data indicators in the set of pre-treatment covariates. In total, we identify 72 pre-treatment covariates of interest.
There is a lack of overlap in the estimated propensity scores in this data and there are many propensity scores close to 1 (supplementary Figure 9). Due to this lack of overlap, there is no solution to Problem (3) and it is difficult to achieve sufficient covariate balance. We estimate the ATE with IPW, minimal weights, and our proposed design- and model-based estimators that use Algorithm (1) to select $g(\bm{x})$. Since rare binary covariates tend to be challenging to exactly balance, we also compute our proposed estimators when initializing Algorithm (1) with the set of rare binary covariates with outcomes that occur for < 5% of observations. We also estimate the ATO using IPW and balancing weights (i.e., all covariates in the set $g(\bm{x})$). We then compare the covariate balance, treatment effect estimates, and Wald-type bootstrapped 95% confidence intervals (CIs) between all methods.
The treatment effect estimates and 95% confidence CIs are similar for all the estimators except the IPW estimator, which corresponds to a more negative estimate and a larger confidence interval (Figure 3). However, all estimates and 95% CIs indicate that IAC non-significantly reduces 28 day mortality by a small percentage. The fact that the minimal weights, ATO, and our proposed estimates are similar likely indicates that there is limited treatment effect heterogeneity. Thus, the main difference between these estimators for this application are in terms of covariate balance and target population. Figure 3 shows the weighted treated and control sample for three pre-treatment covariates in comparison to the sample mean for those covariates (the black horizontal line). IPW and minimal weights do not exactly balance any of the covariates, with corresponding average standardized mean differences (SMD) of $0.161$ and $0.008$, respectively. In contrast, all other weights exactly balance the treated and control distributions. Our proposed design-based weights where Algorithm 1 is initiated with rare binary covariates corresponds to the smallest $g(\bm{x})$ set. In Figure 3, these weights balance both diastolic pressure and the SOFA score to the target population, which is not true of any of the other estimators. Supplementary Figures 10-12 show the SMD for all covariates 1) between treated and control groups; and 2) between treated/control groups and the target population. While these estimators yield similar estimates and 95% confidence intervals for this application, our proposed estimators tend to have improved covariate balance and/or correspond to a minimally modified target population.
Figure 3: ATE estimates and 95% confidence intervals of the effect of IAC on 28-day mortality (bottom figure) and weighted covariate balance of three covariates (three top figures) for eight different weighting/estimation methods (indicated by color). For the top three figures, the horizontal line indicates the sample mean of the covariates in the target population and vertical lines indicate the extent of covariate imbalance for each method.
Transporting “health care hotspotting” RCT to a Midwestern U.S. academic health center population
A recent intervention of popular interest, termed ‘health care hotspotting’, aimed to improve healthcare delivery for individuals who interact substantially with the healthcare system. Despite the promising nature of this intervention, a randomized controlled trial (RCT) in Camden, NJ found a null effect of this intervention on readmission rates Finkelstein et al., (2020). However, a re-anlaysis of this trial by Yang et al., (2023) found a significant effect of this intervention in the population of individuals who had a high probability of engagement with the randomized intervention. Since there is evidence that the treatment effect varies by population, we are interested in determining the effect of this intervention in a Midwestern U.S. academic health center population.
In the RCT, individuals were randomized to standard of care or an intensive, targeted follow-up after admission. The outcome of interest is a binary indicator of readmission within 30 days. We include all pre-treatment covariates that are in common between both datasets, which includes demographics and hospital stay characteristics, for a total of 21 covariates. We perform a complete case analysis (< 30 missing observations) with a total of 781 trial population observations and 1305 target population observations. There is minimal overlap in the estimated probability of being in the target population (supplementary Figure 13). However, the primary driver of this lack of overlap is race; the trial populations is 84.9% non-white while the target population is only 13.6% non-white. When race is removed from the model predicting the probability of being in the target population, there is substantial overlap (supplementary Figure 13). Thus, we implement our proposed design-based estimator and our weighted estimator when $g(\bm{x})$ a priori only includes race. We compare this estimator to the normalized inverse odds weighted (IOW) estimator (Dahabreh et al.,, 2020), the inverse odds weighted estimator where population probabilities less than $0.1$ and more than $0.9$ are trimmed (Crump et al.,, 2009), and the standard stable balancing weights estimator. We report covariate balance, transported treatment effect estimates, and and Wald-type bootstrapped 95% CIs for these methods.
The trimmed estimator has the largest treatment effect estimate, but also the largest 95% CI, likely due to the small sample size after trimming (Figure 4). The IOW and balancing weights estimates and 95% CIs indicate a non-significant reduction in 30-day readmission; however, there was no solution to the direct balancing weights optimization problem for many of the bootstrapped datasets such that a true 95% bootstrapped CI cannot be derived. While the 95% CIs for both of our proposed estimators do not include zero, the interval is substantially smaller and the estimates indicate that the intervention may actually increase hospital readmission rates in the target population. Supplementary Figures 14-16 show the SMD for all covariates 1) between treated and control groups; and 2) between treated/control groups and the target population. Both the IOW and trimmed estimators achieve poor covariate balance with average SMDs of $0.140$ and $0.382$. Balancing weights and our proposed weights achieve exact covariate balance; however, our proposed weights do modify the sample mean of the race covariate such that the estimate applies to a population that is approximately 80% non-white. Therefore, our proposed estimator indicates that the intervention may in fact be detrimental for 30-day readmission rates for this slightly modified target population and for original target population when race is not a treatment effect modifier.
Figure 4: Transported treatment effect estimates and 95% confidence intervals of the effect of ‘health care hotspotting’ on 30-day readmission rates for a Midwestern healthcare center (bottom figure) and weighted covariate balance of two covariates (two top figures) for for different weighting/estimation methods (indicated by color). For the top two figures, the horizontal line indicates the sample mean of the covariates in the target population and vertical lines indicate the extent of covariate imbalance for each method.
Discussion
Positivity violations create substantial challenges when estimating causal effects with observational data. These challenges necessitate researchers to confront trade-offs between estimator bias, variance, and maintaining the original target population. We have proposed a novel weighting procedure that achieves reduced estimator bias and variance through modifying the target population for only a subset of covariates. In doing so, our proposed weights can be derived under an relaxed positivity assumption. We have shown that our proposed estimator 1) is consistent for the original estimand when either the implied propensity score model is correct or the set of treatment effect modifiers is properly specified; and 2) is consistent for a slightly modified estimand that is simple to characterize when these conditions do not hold. Furthermore, we have shown that our proposed estimator achieves reduced asymptotic variance under the relaxed positivity assumption. We have proposed an algorithm for identifying the set of covariates not balanced to the target population; the corresponding design- and model-based estimators perform well across applications to synthetic data, EHR data, and when transporting RCT effects. Thus, our methods allow analysts to derive weights that result in estimators with minimal bias and variance with respect to a meaningful and interpretable target population.
While we focus on estimating the ATE, our proposed weighting procedure can be readily extended to other causal effect estimation problems as discussed in Section 5 and supplementary Section 3. Future work includes exploring the theoretical properties of these extensions. Our proposed extension to distributional balancing weights may be especially promising as these weights do not require strong parametric assumptions on the outcome models. In addition, there is the potential for combining the intuition behind our proposed method and modified treatment policies to obtain estimators with reduced bias and variance that correspond to both a treatment and population of interest. While our weighting procedure does perform well when there are positivity violations with higher dimensional covariates (e.g, $p=100$), there is a need for additional methods that address the combined challenges of variable selection (in the context of unmeasured confounding) and positivity violations in $p>>n$ scenarios. Further, while the infeasibility of Problem (3) is a clear indication that our proposed method or minimal weights need to be used to derive weights, it is generally unclear when estimator performance may be impacted by the overlap levels for a given research problem. Thus, future work includes developing measures of overlap that are directly related the impact of overlap on estimator MSE.
Supplementary Materials for Partially Retargeted Balancing Weights for Causal Effect Estimation Under Positivity Violations
Technical results and proofs
For all proofs, we notate $||\cdot||$ as the $L2$ norm.
Theorem 1
Proof.
We rewrite the problem in matrix notation:
$\text{argmin}_{w_{i}}$ $\>\>\sum_{i=1}^{n}h(s_{i})$
subject to $\>\>Q_{2K+L\times 2n}s_{2n\times 1}=0_{2K+L\times 1}$
where
$s_{2n\times 1}$ $=\begin{bmatrix}(1-Z_{i}w_{i})_{n\times 1}\\ (1-(1-Z_{i})w_{i})_{n\times 1}\end{bmatrix}$
$Q_{2K+L\times 2n}$ $=\begin{bmatrix}c_{1}(\bm{X}_{1})&c_{1}(\bm{X}_{2})&\cdots&c_{1}(\bm{X}_{n})&0&0&\cdots&0\\ c_{K}(\bm{X}_{1})&c_{K}(\bm{X}_{2})&\cdots&c_{K}(\bm{X}_{n})&0&0&\cdots&0\\ 0&0&\cdots&0&c_{1}(\bm{X}_{1})&c_{1}(\bm{X}_{2})&\cdots&c_{1}(\bm{X}_{n})&\\ 0&0&\cdots&0&c_{K}(\bm{X}_{1})&c_{K}(\bm{X}_{2})&\cdots&c_{K}(\bm{X}_{n})&\\ g_{1}(\bm{X}_{1})&g_{1}(\bm{X}_{2})&\cdots&g_{1}(\bm{X}_{n})&-g_{1}(\bm{X}_{1})&-g_{1}(\bm{X}_{2})&\cdots&-g_{1}(\bm{X}_{n})\\ g_{L}(\bm{X}_{1})&g_{L}(\bm{X}_{2})&\cdots&g_{L}(\bm{X}_{n})&-g_{L}(\bm{X}_{1})&-g_{L}(\bm{X}_{2})&\cdots&-g_{L}(\bm{X}_{n})\\ \end{bmatrix}$
This problem is now in the form of Tseng and Bertsekas, (1991). The dual of this problem is,
$\text{minimize}_{\lambda}$ $\>\>\>q(\lambda)$
subject to $\>\>\>\text{no constraint on}\>\>\lambda$
where $q(\lambda)=\sum_{j=1}^{2n}h^{*}_{j}(Q_{j}^{T}\lambda)$ and $h^{*}_{j}(t)=\sup_{s_{j}}\{ts_{j}-h(s_{j})\}$. Consider $j<n+1$. Then we have that,
$h^{*}_{j}(t)$ $=\sup_{s_{j}}\{ts_{j}-h(s_{j})\}$
$=\sup_{w_{j}}\{t-tZ_{j}w_{j}-h(1-Z_{j}w_{j})\}$
$=\sup_{w_{j}}\{t-tZ_{j}w_{j}-Z_{j}h(1-w_{j})-(1-Z_{j})h(1)\}$
$=t-tZ_{j}w_{j}^{*}-Z_{j}h(1-w_{j}^{*})-(1-Z_{j})h(1)$
Next we have that,
$h^{*}_{j+n}(t)$ $=\sup_{s_{j+n}}\{ts_{j+n}-h(s_{j+n})\}$
$=\sup_{w_{j}}\{t-t(1-Z_{j})w_{j}-h[1-(1-Z_{j})w_{j}]\}$
$=t-t(1-Z_{j})w_{j}^{*}-(1-Z_{j})h(1-w_{j}^{*})-Z_{j}h(1)$
where $w_{j}^{*}$ satisfies the first order conditions:
$-tZ_{j}+Z_{j}h^{\prime}(1-w^{*}_{j})$ $=0$
$-t(1-Z_{j})+(1-Z_{j})h(1-w_{j}^{*})$ $=0$
  $\implies\>\>1-(h^{\prime})^{-1}(t)$ $=w^{*}_{j}$
Then we have that,
$h^{*}_{j}(t)$ $=t-tZ_{j}(1-(h^{\prime})^{-1}(t))-Z_{j}h(1-1+(h^{\prime})^{-1}(t))-(1-Z_{j})h(1)$
$=t-tZ_{j}\{1-(h^{\prime})^{-1}(t)\}-Z_{j}h\{(h^{\prime})^{-1}(t)\}-(1-Z_{j})h(1)$
$=-Z_{j}\{t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(-1)\}-h(1)+t$
$h^{*}_{j+n}(t)$ $=t-t(1-Z_{j})\{1-(h^{\prime})^{-1}(t)\}-(1-Z_{j})h\{1-1+(h^{\prime})^{-1}(t)\}-Z_{j}h(1)$
$=t-t(1-Z_{j})\{1-(h^{\prime})^{-1}(t)\}-(1-Z_{j})h\{(h^{\prime})^{-1}(t)\}-Z_{j}h(1)$
$=-(1-Z_{j})\{t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(1)\}-h(1)+t$
Let $\rho(t)=t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(1)$. This gives,
$h^{*}_{j}(t)$ $=-Z_{j}\rho(t)-h(1)+t$
$h^{*}_{j+n}(t)$ $=-(1-Z_{j})\rho(t)-h(1)+t$
Now, note that
$\rho^{\prime}(t)$ $=1-(h^{\prime})^{-1}(t)-t\{(h^{\prime})^{-1}(t)\}^{\prime}+h^{\prime}[(h^{\prime})^{-1}(t)]\times\{(h^{\prime})^{-1}(t)\}^{\prime}$
$=1-(h^{\prime})^{-1}(t)-t\{(h^{\prime})^{-1}(t)\}^{\prime}+t\{(h^{\prime})^{-1}(t)\}^{\prime}$
$=1-(h^{\prime})^{-1}(t)$
$\implies\rho^{\prime}(t)$ $=w_{j}^{*}$
Thus, the dual formulation becomes
$\text{minimize}_{\lambda}$ $\>\>\>q(\lambda)$
subject to $\>\>\>\text{no constraint on}\>\>\lambda$
where
$q(\lambda)$ $=\frac{1}{n}\sum_{j=1}^{n}-Z_{j}\rho(Q_{j}^{T}\lambda)-(1-Z_{j})\rho(Q_{j+n}^{T}\lambda)+Q_{j}^{T}\lambda+Q_{j+n}^{T}\lambda-2nh(1)$
$=\frac{1}{n}\sum_{j=1}^{n}-Z_{j}\rho(Q_{j}^{T}\lambda)-(1-Z_{j})\rho(Q_{j+n}^{T}\lambda)+Q_{j}^{T}\lambda+Q_{j+n}^{T}\lambda$
Let $B^{+}(\bm{X}_{j})=\{c_{1}(\bm{X}_{j}),\ldots,c_{K}(\bm{X}_{j}),0_{K},g_{1}(\bm{X}_{j}),\ldots,g_{L}(\bm{X}_{j})\}$
and $B^{-}(\bm{X}_{j})=\{0_{K},c_{1}(\bm{X}_{j}),\ldots,c_{K}(\bm{X}_{j}),-g_{1}(\bm{X}_{j}),\ldots,-g_{L}(\bm{X}_{j})\}$. Then, the primal solution $w_{j}^{*}$ satisfies
$w^{*}_{j}$ $=\rho^{\prime}\{B^{+}(\bm{X}_{j})^{T}\hat{\lambda}\}\>\>\text{for}\>\>Z_{j}=1\>\>;\>\>j=1,\ldots n$
$w^{*}_{j}$ $=\rho^{\prime}\{B^{-}(\bm{X}_{j})^{T}\hat{\lambda}\}\>\>\text{for}\>\>Z_{j}=0\>\>;\>\>j=1,\ldots n$
where $\hat{\lambda}$ is the solution to the dual problem. Let $\lambda=\begin{bmatrix}\bm{\alpha}_{0_{K\times 1}}\\ \bm{\alpha}_{1_{K\times 1}}\\ \bm{\gamma}_{L\times 1}\end{bmatrix}$ and let $c(\bm{X}_{j})=\{c_{1}(\bm{X}_{j}),\ldots c_{K}(\bm{X}_{j})\}$ and $g(\bm{X}_{j})=\{g_{1}(\bm{X}_{j}),\ldots,g_{L}(\bm{X}_{j})\}$. Then the dual formulation simplifies to:
$\underset{\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}}{\textnormal{minimize}}$ $\>\>\>q(\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma})$
subject to $\>\>\>\text{no constraint on}\>\>\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}$
$\text{where}\>\>q(\alpha_{0},\alpha_{1},\gamma)$ $=\frac{1}{n}\sum_{i=1}^{n}-Z_{i}\rho\{\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X}_{i})\}-(1-Z_{i})\rho\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X}_{i})\}$
$+\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X}_{i})+\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X}_{i})$
$=\frac{1}{n}\sum_{i=1}^{n}-Z_{i}\rho\{\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X}_{i})\}-(1-Z_{i})\rho\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X}_{i})\}$
$+\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\alpha}_{0}^{T}c(\bm{X}_{i})$
and the primal solution $w_{i}^{*}=Z_{i}w^{*}_{1i}+(1-Z_{i})w^{*}_{0i}$ satisfies $w^{*}_{1i}=\rho^{\prime}\{\hat{\bm{\alpha}}_{1}^{T}c(\bm{X}_{i})+\hat{\bm{\gamma}}^{T}g(\bm{X}_{i})\}\>\>\text{and}\>\>w^{*}_{0i}=\rho^{\prime}\{\hat{\bm{\alpha}}_{0}^{T}c(\bm{X}_{i})-\hat{\bm{\gamma}}^{T}g(\bm{X}_{i})\}$ where $\hat{\bm{\alpha}}_{0},\hat{\bm{\alpha}}_{1},\hat{\bm{\gamma}}$ is the solution to the dual problem. ∎
Proposition 2
Proof.
[…]
Let $R_{i}(\bm{x}^{*})=i=1,\ldots 3^{p}$ be the $3^{p}$ boxes centered at $\bm{x}^{*}+\frac{3}{2}\epsilon b$ where $b\in\mathbb{R}^{p}$ contains entries of 0,1,10,-1,1. Then, $P(B_{\epsilon}(\bm{x}^{*})\in\text{CH}\{\bm{X}_{i}\}_{Z_{i}=1},B_{\epsilon}(\bm{x}^{*})\in\text{CH}\{\bm{X}_{i}\}_{Z_{i}=0})\geq P(\exists\bm{X}_{i}\in R_{i}(\bm{x}^{*}),Z_{i}=0,\exists\bm{X}_{j}\in R_{j}(\bm{x}^{*}),Z_{j}=1,i,j=1,\ldots,3^{p})$.
Assume that for some $\bm{x}^{*}$, for all $\bm{x}\in B_{2\epsilon}(\bm{x}^{*}),\>\>\>0<P(Z_{i}|\bm{X}_{i}=\bm{x})<1$. This assumption implies that $\rho_{1}=\min_{i}P(\bm{X}\in R_{i}(\bm{x}^{*})|Z=1)>0$ and $\rho_{0}=\min_{i}P(\bm{X}\in R_{i}(\bm{x}^{*})|Z=0)>0$. Then we have the following,
$P(\exists\bm{X}_{i}\in R_{i}(\bm{x}^{*}),Z_{i}=0,$ $\exists\bm{X}_{j}\in R_{j}(\bm{x}^{*}),Z_{j}=1,i,j=1,\ldots,3^{p})$
$=P(\exists\bm{X}_{i}\in R_{i}(\bm{x}^{*}),Z_{i}=0,i=1,\ldots,3^{p})$
$\times P(\exists\bm{X}_{j}\in R_{j}(\bm{x}^{*}),Z_{j}=1,j=1,\ldots,3^{p})$
$\geq\{1-\sum_{i=1}^{3^{p}}P(X\not\in R_{i}(\bm{x}^{*})|Z=0)^{n}\}$
$\times\{1-\sum_{i=1}^{3^{p}}P(X\not\in R_{j}(\bm{x}^{*})|Z=1)^{n}\}$
$=\{1-3^{p}(1-\rho_{0})^{n}\}\{1-3^{p}(1-\rho_{1})^{n}\}$
$\lim_{n\to\infty}$ $\to 1$
where the first equality holds because each $(\bm{X}_{i},Z_{i})$ is an iid sample. We see that letting $\epsilon\to 0$, the only assumption required is the existence of some $\bm{x}^{*}$ such that $0<P(Z_{i}|\bm{X}_{i}=\bm{x}^{*})<1$.
Proposition 5
Proposition 5.
Suppose Assumptions Assumption 1: Strong ignorability, either Assumption 2: Positivity or Assumption 2: Relaxed positivity, and Assumption 3: Balance function regularity conditions hold and $\rho$ has a continuous first derivative. Let $\bm{\theta}=(\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}).$ Then,
[…]
ii) $w^{*}_{1}(\bm{X})=\rho^{\prime}\{\hat{\bm{\alpha}}^{T}_{1}c(\bm{X})+\hat{\bm{\gamma}}^{T}g(\bm{X})\}\to_{p}\tilde{w}_{1}(\bm{X})=\rho^{\prime}\{\tilde{\bm{\alpha}}^{T}_{1}c(\bm{X})+\tilde{\bm{\gamma}}^{T}g(\bm{X})\}$ and $w^{*}_{0}(\bm{X})=\rho^{\prime}\{\hat{\bm{\alpha}}_{0}^{T}c(\bm{X})+-\hat{\bm{\gamma}}^{T}g(\bm{X})\}\to_{p}\tilde{w}_{0}(\bm{X})=\rho^{\prime}\{\tilde{\bm{\alpha}}^{T}_{0}c(\bm{X})-\tilde{\bm{\gamma}}^{T}g(\bm{X})\}.$
We use similar proof techniques to the proof of proposition 1 in Källberg and Waernbaum, (2023). First, we prove the following lemma.
Lemma 6.
[…]
Proof.
[…]
$H_{k_{1},k_{2}}(\bm{\theta})$ $=E[-e(\bm{X})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X})+\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})c_{k_{2}}(\bm{X})$
$-\{1-e(\bm{X})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X})-\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})c_{k_{2}}(\bm{X})],$
$H_{l_{1},l_{2}}(\bm{\theta})$ $=E[-e(\bm{X})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X})+\bm{\gamma}^{T}g(\bm{X})\}g_{l_{1}}(\bm{X})g_{l_{2}}(\bm{X})$
$-\{1-e(\bm{X})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X})\}g_{l_{1}}(\bm{X})g_{l_{2}}(\bm{X})],$
$H_{k_{1},l_{1}}(\bm{\theta})$ $=E[-e(\bm{X})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})g_{l_{1}}(\bm{X})$
$-\{1-e(\bm{X})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X})\}c_{k_{1}}(\bm{X})g_{l_{1}}(\bm{X})].$
We can construct an estimator for $H(\bm{\theta})$, $\hat{H}(\bm{\theta})$ using the sample equivalents of $H_{k_{1},k_{2}}(\bm{\theta}),H_{l_{1},l_{2}}(\bm{\theta}),$ and $H_{k_{1},l_{1}}(\bm{\theta})$. Let $c_{i}=-e(\bm{X}_{i})\rho^{\prime\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X}_{i})+\bm{\gamma}^{T}g(\bm{X}_{i})\}+-\{1-e(\bm{X_{i}})\}\rho^{\prime\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X}_{i})-\bm{\gamma}^{T}g(\bm{X_{i}})\}$. By the strict concavity of $\rho$, $c_{i}>0$. Let $v^{*}=[\sqrt{c_{1}},\ldots,\sqrt{c_{n}}]$. Let $v$ be any vector of length $n$. Then $v^{T}\hat{H}(\bm{\theta})v=v^{T}\odot v^{*^{T}}B(\bm{X})^{T}B(\bm{X})v\odot v^{*}>0$ since $B(\bm{X})$ is full rank by Assumption Assumption 3: Balance function regularity conditions. Then $\hat{H}$ is positive definite and since $E[\hat{H}(\bm{\theta})]=H(\bm{\theta})$, $\hat{H}(\bm{\theta})\to_{p}H(\bm{\theta})$, $H(\bm{\theta})$ is positive definite and $l(\bm{\theta})$ is strictly convex as desired. ∎
Now we will prove Proposition 5.
Proof.
[…]
Then since $\rho^{\prime}$ is continuous, by the continuous mapping theorem we have that $w^{*}_{1}(\bm{X})=\rho^{\prime}\{\hat{\bm{\alpha}}^{T}_{1}c(\bm{X})+\hat{\bm{\gamma}}^{T}g(\bm{X})\}\to_{p}\tilde{w}_{1}(\bm{X})=\rho^{\prime}\{\tilde{\bm{\alpha}}^{T}_{1}c(\bm{X})+\tilde{\bm{\gamma}}^{T}g(\bm{X})\}$ and $w^{*}_{0}(\bm{X})=\rho^{\prime}\{\hat{\bm{\alpha}}_{0}^{T}c(\bm{X})-\hat{\bm{\gamma}}^{T}g(\bm{X})\}\to_{p}\tilde{w}_{0}(\bm{X})=\rho^{\prime}\{\tilde{\bm{\alpha}}^{T}_{0}c(\bm{X})-\tilde{\bm{\gamma}}^{T}g(\bm{X})\}.$
Theorem 3
Proof.
Assume that the true models are $\{e(\bm{x})\}^{-1}=\rho^{\prime}\{\tilde{\bm{\alpha}}_{1}^{T}c(\bm{x})+\tilde{\bm{\gamma}}^{T}g(\bm{x})\}$ and $\{1-e(\bm{x})\}^{-1}=\rho^{\prime}\{\tilde{\bm{\alpha}}_{0}^{T}c(\bm{x})-\tilde{\bm{\gamma}}^{T}g(\bm{x})\}$. Then given Assumptions Assumption 1: Strong ignorability and Assumption 2: Positivity,
$\tau$ $=E[\frac{Z_{i}Y_{i}}{e(\bm{X}_{i})}]-E[\frac{(1-Z_{i})Y_{i}}{1-e(\bm{X}_{i})}]$
$=E[\rho^{\prime}\{\tilde{\bm{\alpha}}_{1}^{T}c(\bm{x})+\tilde{\bm{\gamma}}^{T}g(\bm{x})\}Z_{i}Y_{i}]-E[\rho^{\prime}\{\tilde{\bm{\alpha}}_{0}^{T}c(\bm{x})-\tilde{\bm{\gamma}}^{T}g(\bm{x})\}(1-Z_{i})Y_{i}].$
Further,
$\tau_{w^{*}}$ $=\sum_{i=1}^{n}w^{*}(\bm{X})Z_{i}Y_{i}-\sum_{i=1}^{n}w^{*}(\bm{X})(1-Z_{i})Y_{i}$
$=\sum_{i=1}^{n}\rho^{\prime}\{\hat{\bm{\alpha}}^{T}_{1}c(\bm{X})+\hat{\bm{\gamma}}^{T}g(\bm{X})\}Z_{i}Y_{i}-\sum_{i=1}^{n}\rho^{\prime}\{\hat{\bm{\alpha}}_{0}^{T}c(\bm{X})-\hat{\bm{\gamma}}^{T}g(\bm{X})\}(1-Z_{i})Y_{i}$
Then by Proposition 5 and the law of large numbers for averages with estimated parameters (Boos and Stefanski, (2013) Theorem 7.3),
$\sum_{i=1}^{n}\rho^{\prime}\{\hat{\bm{\alpha}}^{T}_{1}c(\bm{X})+\hat{\bm{\gamma}}^{T}g(\bm{X})\}Z_{i}Y_{i}$ $\to_{p}E[\rho^{\prime}\{\tilde{\bm{\alpha}}_{1}^{T}c(\bm{x})+\tilde{\bm{\gamma}}^{T}g(\bm{x})\}Z_{i}Y_{i}],$
$\sum_{i=1}^{n}\rho^{\prime}\{\hat{\bm{\alpha}}_{0}^{T}c(\bm{X})-\hat{\bm{\gamma}}^{T}g(\bm{X})\}(1-Z_{i})Y_{i}$ $\to_{p}E[\rho^{\prime}\{\tilde{\bm{\alpha}}_{0}^{T}c(\bm{x})-\tilde{\bm{\gamma}}^{T}g(\bm{x})\}(1-Z_{i})Y_{i}].$
It follows directly that $\tau_{w^{*}}\to_{p}\tau.$
Proof.
Assume $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}\in\mathbb{R}^{L}$. We will use a similar proof technique by Zhao and Percival, (2017) to show that augmenting the our proposed estimator with estimated outcome regressions does not change the estimator. Let $\hat{\mu}_{0}(\bm{x})$, $\hat{\mu}_{1}(\bm{x})$, and $\hat{e}(\bm{x})$ be outcome regression and propensity score estimates. Then the standard doubly robust estimator for the ATE is,
$\hat{\tau}_{DR}=\frac{1}{n_{1}}\sum_{Z_{i}=1}\left[\frac{Y_{i}-\hat{\mu}_{1}(\bm{X}_{i})}{\hat{e}(\bm{X}_{i})}-\hat{\mu}_{1}(\bm{X}_{i})\right]-\frac{1}{n_{0}}\sum_{Z_{i}=1}\left[\frac{Y_{i}-\hat{\mu}_{0}(\bm{X}_{i})}{1-\hat{e}(\bm{X}_{i})}-\hat{\mu}_{0}(\bm{X}_{i})\right]$
Now consider using $w^{*}$ to estimate the propensity score. Then,
$\hat{\tau}_{DR}-\hat{\tau}_{w^{*}}$ $=-\sum_{Z_{i}=1}w^{*}(\bm{X}_{i})\hat{\mu}_{1}(\bm{X}_{i})+\frac{1}{n}\sum_{i}\hat{\mu}_{1}(\bm{X}_{i})+\sum_{Z_{i}=0}w(\bm{X}_{i})^{*}\hat{\mu}_{0}(\bm{X}_{i})-\frac{1}{n}\sum_{i}\hat{\mu}_{0}(\bm{X}_{i})$
$=-\sum_{Z_{i}=1}w^{*}(\bm{X}_{i})[\hat{\bm{\beta}}_{1}^{T}c(\bm{X}_{i})+\hat{\bm{\lambda}}^{T}g(\bm{X})_{i}]+\frac{1}{n}\sum_{i}\hat{\bm{\beta}}_{1}^{T}c(\bm{X}_{i})+\hat{\bm{\lambda}}^{T}g(\bm{X})_{i}$
$+\sum_{Z_{i}=0}w^{*}(\bm{X}_{i})[\hat{\bm{\beta}}_{0}^{T}c(\bm{X}_{i})+\hat{\bm{\lambda}}^{T}g(\bm{X})_{i}]-\frac{1}{n}\sum_{i}\hat{\bm{\beta}}_{0}^{T}c(\bm{X}_{i})+\hat{\bm{\lambda}}^{T}g(\bm{X})_{i}$
$=\hat{\bm{\beta}}_{1}^{T}[\sum_{i}c(\bm{X}_{i})-\sum_{i}w^{*}(\bm{X}_{i})Z_{i}c(\bm{X}_{i})]$
$-\hat{\bm{\beta}}_{0}^{T}[\sum_{i}c(\bm{X}_{i})-\sum_{i}w^{*}(\bm{X}_{i})(1-Z_{i})c(\bm{X}_{i})]$
$+\hat{\bm{\lambda}}^{T}[\sum_{i}w^{*}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})-\sum_{i}w^{*}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})]$
$=0,$
and $\hat{\tau}_{DR}=\hat{\tau}_{w^{*}}$. Since $\hat{\tau}_{DR}\to_{p}\tau$ we also have $\hat{\tau}_{w^{*}}\to_{p}\tau$. ∎
Corollary 3.1
Proof.
Assume $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}_{1}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}_{0}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}_{z}\in\mathbb{R}^{L}$. Then by Proposition 3 and the law of large numbers for averages with estimated parameters (Boos and Stefanski, (2013) Theorem 7.3)
$\frac{1}{n}\sum_{i=1}^{n}w^{*}(\bm{X})Z_{i}Y_{i}$ $\to_{p}E[\tilde{w}(\bm{X}_{i})Z_{i}Y_{i}]$
$=E[\tilde{w}(\bm{X}_{i})Z_{i}\{\bm{\beta}_{1}^{T}c(\bm{X}_{i})+\bm{\lambda}_{1}^{T}g(\bm{X}_{i})\}]$
$=E[\bm{\beta}_{1}^{T}c(\bm{X}_{i})+\tilde{w}(\bm{X}_{i})Z_{i}\bm{\lambda}_{1}^{T}g(\bm{X}_{i})]$
$=E[\bm{\beta}_{1}^{T}c(\bm{X}_{i})+\bm{\lambda}_{1}^{T}\tilde{w}(\bm{X}_{i})e(\bm{X}_{i})g(\bm{X}_{i})]$
$=E[E[Y_{i}(1)|c(\bm{X}_{i}),\tilde{w}(\bm{X}_{i})e(\bm{X}_{i})g(\bm{X}_{i})]]$
$=E[E[Y_{i}|Z_{i}=1,c(\bm{X}_{i}),\tilde{w}(\bm{X}_{i})e(\bm{X}_{i})g(\bm{X}_{i})]].$
The second equality holds by $\tilde{w}$ satisfying the first order conditions of the population version of Equation (6). Similarly, $\frac{1}{n}\sum_{i=1}^{n}w^{*}(\bm{X})(1-Z_{i})Y_{i}\to E[E[Y_{i}|Z_{i}=0,c(\bm{X}_{i}),\tilde{w}(\bm{X}_{i})\{1-e(\bm{X}_{i})\}g(\bm{X}_{i})]]$. Then $\hat{\tau}_{w^{*}}\to\tau_{g,\tilde{w}}$ as desired. ∎
Theorem 4
Proof.
We can decompose $\hat{\tau}_{w^{*}}-\tau$ into the following:
$\hat{\tau}_{w^{*}}-\tau$ $=\frac{1}{n}\sum_{i=1}^{n}w^{*}(\bm{X}_{i})Z_{i}Y_{i}-w^{*}(\bm{X}_{i})(1-Z_{i})Y_{i}-\tau$
$=\frac{1}{n}\sum_{i=1}^{n}w^{*}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}+\frac{1}{n}\sum_{i=1}^{n}(w^{*}_{1}(\bm{X}_{i})Z_{i}-1)\mu_{1}(\bm{X}_{i})$
$-\frac{1}{n}\sum_{i=1}^{n}w^{*}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}-\mu_{0}(\bm{X}_{i})\}+\frac{1}{n}\sum_{i=1}^{n}\{w^{*}_{0}(\bm{X}_{i})(1-Z_{i})-1\}\mu_{0}(\bm{X}_{i})$
$+\frac{1}{n}\sum_{i=1}^{n}\tau(\bm{X}_{i})-\tau$
$=\frac{1}{n}\sum_{i=1}^{n}S_{i}+R_{1}+R_{2}$
where,
$S_{i}$ $=\frac{Z_{i}}{e(\bm{X}_{i})}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}-\frac{1-Z_{i}}{1-e(\bm{X}_{i})}\{Y_{i}-\mu_{0}(\bm{X}_{i})\}+\{\tau(\bm{X}_{i})-\tau\}$
$R_{1}$ $=\frac{1}{n}\sum_{i=1}^{n}(w^{*}_{1}(\bm{X}_{i})-\frac{1}{e(\bm{X}_{i})})Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}-(w^{*}_{0}(\bm{X}_{i})-\frac{1}{1-e(\bm{X}_{i})})(1-Z_{i})\{Y_{i}-\mu_{0}(\bm{X}_{i})\}$
$R_{2}$ $=\frac{1}{n}\sum_{i=1}^{n}(w_{1}^{*}(\bm{X}_{i})Z_{i}-1)\mu_{1}(\bm{X}_{i})\}]-\{w_{0}^{*}(\bm{X}_{i})(1-Z_{i})-1\}\mu_{0}(\bm{X}_{i})$
$=\frac{1}{n}\sum_{i=1}^{n}(w^{*}_{1}(\bm{X}_{i})Z_{i}-1)\{\bm{\beta}_{1}^{T}c(\bm{X}_{i})+\bm{\lambda}^{T}g(\bm{X}_{i})\}-\{w^{*}_{0}(\bm{X}_{i})(1-Z_{i})-1\}\{\bm{\beta}_{1}^{T}c(\bm{X}_{i})+\bm{\lambda}^{T}g(\bm{X}_{i})\}$
$=\frac{1}{n}\sum_{i=1}^{n}(w^{*}_{1}(\bm{X}_{i})Z_{i}-1)\bm{\beta}_{1}^{T}c(\bm{X}_{i})-\sum_{i=1}^{n}\{w_{0}^{*}(\bm{X}_{i})(1-Z_{i})-1\}\bm{\beta}_{0}^{T}c(\bm{X}_{i})$
$+\sum_{i=1}^{n}\{w_{1}^{*}(\bm{X}_{i})Z_{i}-w_{0}^{*}(\bm{X}_{i})(1-Z_{i})\}\bm{\lambda}^{T}g(\bm{X}_{i})$
Note that $R_{2}=0$ by the constraints in the optimization problem. Then, since $S_{i}$ takes the same form as the efficient score for the ATE in Hahn, (1998), $\hat{\tau}_{w^{*}}$ is asymptotically normal and semiparametrically efficient as long as $R_{1}$ is $o_{p}(n^{-1/2}).$
We will focus on the first component of $R_{1}$,
$|R_{1}^{1}|$ $=|\frac{1}{n}\sum_{i=1}^{n}(w^{*}_{1}(\bm{X}_{i})-\frac{1}{e(\bm{X}_{i})})Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}|$
$=|\int(w^{*}_{1}(\bm{X}_{i})-\frac{1}{e(\bm{X}_{i})})\{Z_{i}Y_{i}-Z_{i}\mu_{1}(\bm{X}_{i})\}dF_{n}(\bm{x})|$
[…]
$=o_{p}(1)O_{p}(n^{-1/2})=o_{p}(n^{-1/2})$
where the third inequality follows from Cauchy-Schwartz and the fourth equality follows from Theorem 1 and and the central limit theorem. Similar arguments hold for $R_{1}^{2}$ and we get that
$R_{1}$ $\leq|R_{1}^{1}|+|R_{1}^{2}|$
$=o_{p}(n^{-1/2})+o_{p}(n^{-1/2})=o_{p}(n^{-1/2})$
as desired. ∎
Corollary 4.1
Proof.
When the propensity score model is not correctly specified, $R_{1}$ will no longer go to zero. Then, the variance of $\hat{\tau}_{w^{*}}$ will be $\text{Var}(S_{i}+R_{1_{i}})$ where
$S_{i}+R_{1_{i}}$ $=w_{1}^{*}(\bm{X}_{i})Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}-w_{0}^{*}(\bm{X}_{i})(1-Z_{i})\{Y_{i}-\mu_{0}(\bm{X}_{i})\}+\{\tau(\bm{X}_{i})-\tau\}$
$=w_{1}^{*}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-w_{0}^{*}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}+\{\tau(\bm{X}_{i})-\tau\}$
$=\{w_{1}^{*}(\bm{X}_{i})-\tilde{w}_{1}(\bm{X}_{i})\}Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-\{w_{0}^{*}(\bm{X}_{i})-\tilde{w}_{0}(\bm{X}_{i})\}(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}$
$+\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}+\{\tau(\bm{X}_{i})-\tau\}$
Using similar arguments used for $R_{1}$ above and Proposition 3,
$\{w_{1}^{*}(\bm{X}_{i})-\tilde{w}_{1}(\bm{X}_{i})\}Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-\{w_{0}^{*}(\bm{X}_{i})-\tilde{w}_{0}(\bm{X}_{i})\}(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}=o_{p}(n^{-1/2})$
such that
$\text{Var}(S_{i}+R_{1i})=\text{Var}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}+\{\tau(\bm{X}_{i})-\tau\}]$
Then,
$\text{Var}(S_{i}+R_{1_{i}})$ $=\text{Var}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}]+\text{Var}[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]$
$+\text{Var}\{\tau(\bm{X}_{i})-\tau\}-2\text{Cov}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\},\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]$
$+2\text{Cov}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\},\{\tau(\bm{X}_{i})-\tau\}]$
$-2\text{Cov}[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\},\{\tau(\bm{X}_{i})-\tau\}]$
We have that,
$-2\text{Cov}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)$- $\mu_{1}(\bm{X}_{i})\},\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]$
$=-2E[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}]\times E[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]]$
$=-2E(E[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}])$
$\times E(E[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}|\bm{X}_{i}])$
$=-2E(\tilde{w}_{1}(\bm{X}_{i})E[Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}])$
$\times E(\tilde{w}_{0}(\bm{X}_{i})E[(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}|\bm{X}_{i}])$
(by ignorability) $=-2E[\tilde{w}_{1}(\bm{X}_{i})e(\bm{X}_{i})\{\mu_{1}(\bm{X}_{i})-\mu_{1}(\bm{X}_{i})\}]$
$\times E[\tilde{w}_{0}(\bm{X}_{i})\{1-e(\bm{X}_{i})\}\{\mu_{0}(\bm{X}_{i})-\mu_{0}(\bm{X}_{i})\}]$
$=0\times 0=0$
By a similar argument, $E[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}]E[\{\tau(\bm{X}_{i})-\tau\}]=E[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]E[\{\tau(\bm{X}_{i})-\tau\}]=0$ so it remains to show that $E[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}\{\tau(\bm{X}_{i})-\tau\}]=E[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}\{\tau(\bm{X}_{i})-\tau\}]=0$. Now,
$E[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}\{\tau(\bm{X}_{i})-\tau\}]$ $=E(E[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}\{\tau(\bm{X}_{i})-\tau\}|\bm{X}_{i}])$
$=E(\tilde{w}_{1}(\bm{X}_{i})\{\tau(\bm{X}_{i})-\tau\}E[Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}])$
(by ignorability) $=E[\tilde{w}_{1}(\bm{X}_{i})\{\tau(\bm{X}_{i})-\tau\}e(\bm{X}_{i})\{\mu_{1}(\bm{X}_{i})-\mu_{1}(\bm{X}_{i})\}]$
$=0$
and by a similar argument $E[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}\{\tau(\bm{X}_{i})-\tau\}]=0$ such that all covariance terms are zero. Next,
$\text{Var}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}]$ $=E(\text{Var}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}])$
$+\text{Var}(E[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}])$
(by similar argument to above) $=E(\text{Var}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}])+\text{Var}(0)$
$=E(\tilde{w}_{1}(\bm{X}_{i})^{2}\text{Var}[Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}])$
Now, note that
$\text{Var}\left[Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}\right]$ $=\text{Var}(Z_{i}|\bm{X}_{i})\text{Var}[\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}]$
$+\text{Var}(Z_{i}|\bm{X}_{i})E[\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}]^{2}$
$+\text{Var}[\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}]E[Z_{i}|\bm{X}_{i}]^{2}$
$=e(\bm{X}_{i})\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}+e(\bm{X}_{i})^{2}\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}$
$=e(\bm{X}_{i})(1-e(\bm{X}_{i})+e(\bm{X}_{i}))\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}$
$=e(\bm{X}_{i})\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}$
Then,
$\text{Var}[\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}]$ $=E[\tilde{w}_{1}(\bm{X}_{i})^{2}e(\bm{X}_{i})\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}]$
and by a similar argument,
$\text{Var}[\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}]$ $=E[\tilde{w}_{0}(\bm{X}_{i})^{2}\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(0)|\bm{X}_{i}\}].$
Finally, we have that
$\text{Var}\{\tau(\bm{X}_{i})-\tau\}$ $=E[\{\tau(\bm{X}_{i})-\tau\}^{2}]-E[\tau(\bm{X}_{i})-\tau]^{2}$
$=E[\{\tau(\bm{X}_{i})-\tau\}^{2}]-(E[\tau(\bm{X}_{i})]-\tau)^{2}$
$=E[\{\tau(\bm{X}_{i})-\tau\}^{2}].$
Then, we have that the asymptotic variance of the estimator when the propensity score model is misspecified is,
$E[\tilde{w}_{1}(\bm{X}_{i})^{2}e(\bm{X}_{i})\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}+\tilde{w}_{0}(\bm{X}_{i})^{2}\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(0)|\bm{X}_{i}\}+\{\tau(\bm{X}_{i})-\tau\}^{2}]$
Restrictions on the implied propensity score model
We explore the restrictions of the implied propensity score model commonly used measures of dispersion: 1) entropy weights; and 2) stable balancing weights. The first order conditions in (7) imply the following condition:
$(\rho^{\prime})^{-1}(\{e(\bm{x})\}^{-1})+(\rho^{\prime})^{-1}(\{1-e(\bm{x})\}^{-1})$ $=(\bm{\alpha}_{0}+\bm{\alpha}_{1})^{T}c(\bm{x})$
Without a defining a specific measure of dispersion $D(\cdot)$ (and therefore a specific $\rho(\cdot)$) it is hard to connect this to $Var(Z|\bm{X})$. However, we can explore two common deviance measures (Wang and Zubizarreta,, 2020):
  1. 1.
    Entropy: $D(t)=t\log(t)$ $\implies$ $\rho^{\prime}(t)=e^{-t-1}$ $\implies$ $(\rho^{\prime})^{-1}(t)=-\log(t)-1$
  2. 2.
    Stable balancing weights: $D(t)=t^{2}$ $\implies$ $\rho^{\prime}(x)=-\frac{t}{2}$ $\implies$ $(\rho^{\prime})^{-1}(x)=-2t$
Consider the entropy measure of weight dispersion first. Applying this to the above equations, we get the following
$\log\{e(\bm{x})\}=\bm{\alpha}_{1}^{T}c(\bm{x})+\bm{\gamma}^{T}g(\bm{x})+1$
$\log\{1-e(\bm{x})\}=\bm{\alpha}_{1}^{T}c(\bm{x})-\bm{\gamma}^{T}g(\bm{x})+1$
$\log\{e(\bm{x})\}+\log\{1-e(\bm{x})\}=(\bm{\alpha}_{0}+\bm{\alpha}_{1})^{T}c(\bm{x})+2$
$\log\{e(\bm{x})(1-e(\bm{x})\}=(\bm{\alpha}_{0}+\bm{\alpha}_{1})^{T}c(\bm{x})+2$
$\text{Var}(Z|\bm{X}=\bm{x})=\text{exp}\{(\bm{\alpha}_{0}+\bm{\alpha}_{1})^{T}c(\bm{x})+2\}$
such that $\text{Var}(Z|\bm{X}=\bm{x})$ is modeled as only varying with $c(\bm{x})$. Now, consider the dispersion measure used in the stable balancing weights paper:
$\frac{1}{e(\bm{x})}$ $=-2\bm{\alpha}_{1}^{T}c(\bm{x})-2\bm{\gamma}^{T}g(\bm{x})$
$\frac{1}{1-e(\bm{x})}$ $=-2\bm{\alpha}_{0}^{T}c(\bm{x})+2\bm{\gamma}^{T}g(\bm{x})$
$\frac{1}{e(\bm{x})}+\frac{1}{1-e(\bm{x})}$ $=-2\bm{\alpha}_{1}^{T}c(\bm{x})-2\bm{\alpha}_{0}^{T}c(\bm{x})$
$\frac{1}{e(\bm{x})(1-e(\bm{x}))}$ $=-2\bm{\alpha}_{1}^{T}c(\bm{x})-2\bm{\alpha}_{0}^{T}c(\bm{x})$
$\text{Var}(Z|\bm{X}=\bm{x})$ $=\frac{1}{-2\bm{\alpha}_{1}^{T}c(\bm{x})-2\bm{\alpha}_{0}^{T}c(\bm{x})}$
and again, $\text{Var}(Z|\bm{X}=\bm{x})$ is modeled as only varying with $c(\bm{x})$. While this relationship may not hold for all $D(\cdot)$ functions, we can generally think of this formulation of the direct balancing problem as allowing us to model $E[Z|\bm{X}=\bm{x}]$ with all functions of covariates, but $\text{Var}(Z|\bm{X}=\bm{x})$ with only the set of covariates that are balanced to the population. However, since this in general does not hold for an arbitrary function $g(\bm{x})$, this places substantial restrictions on what forms $g(\bm{x})$ can take for the implied propensity score model to be true. This $g(\bm{x})$ form can be identified for each $D(\cdot)$ by using the first order conditions (7a)-(7c) to solve for $\bm{\gamma}^{T}g(\bm{x})$ as a function of $\bm{\alpha}_{z}^{T}c(\bm{x})$.
Connections to other estimation problems
Multiple treatment groups
Consider the case where there are $M>2$ treatment groups such that $Z_{i}=m$ for $m=0,\ldots,M-1$ and where we want to estimate the average treatment effect between treatments $m_{1},m_{2}=0,\ldots,M-1$, $\tau_{m_{1},m_{2}}=E[\mu_{m_{1}}(\bm{x})-\mu_{m_{2}}(\bm{x})]$. Adapting Problem (5) to derive weights for estimating $\tau_{m_{1},m_{2}}$ simply requires replacing $z=0,1$ with $z=m_{1},m_{2}$ in Problem (5). However, if weights are derived in this way for a series of average treatment effects (e.g., $\tau_{m_{1},m_{2}},\tau_{m_{2},m_{3}},\tau_{m_{1},m_{3}}$ etc.,), each average treatment effect will correspond to a different population for the covariate set $g(\bm{x})$ or even a different set of covariates in $g(\bm{x})$. To maintain the same $g(\bm{x})$ covariate population for average treatment effects corresponding to a set all potential contrasts for $M_{1}\leq M$ treatments, we would include $M_{1}$ constraints balancing $c(\bm{x})$ in treatment groups to the sample population (e.g., constraint (5b)) and $M_{1}-1$ constraints balancing $g(\bm{x})$ between treatment groups (e.g., constraint (5c)). As $M_{1}$ increases, generally the set $g(\bm{x})$ will need to be larger for the existence of a solution. However, for there to be no estimator error, $g(\bm{x})$ must contain no effect modifiers for all treatment groups. Thus, deriving weights for all contrasts separately may allow for a smaller $g(\bm{x})$ set, but this set and corresponding population may vary for each contrast. In contrast, deriving weights for contrasts together will result in the same target population for all average treatment effect estimands but the set $g(\bm{x})$ may need to be large to derive the weights.
Weighted average treatment effects
Consider the estimand $E[h(\bm{X})\tau(\bm{X})]$ for some function $h(\bm{x})$. This as a weighted average treatment effect (WATE) where the population covariate density $f(\bm{x})$ is modified to $h(\bm{x})f(\bm{x})$. The ATO is the WATE when $h(\bm{x})=e(\bm{x})\{1-e(\bm{x})\}$. When adapting our proposed constrained optimization problem to derive weights for estimating the WATE, $\{w_{i}^{WATE}\}$, we can simply replace the right-hand side of constraint (5b) with $\sum_{i=1}^{n}h(\bm{X}_{i})c_{k}(\bm{X}_{i})/\sum_{i=1}^{n}h(\bm{X}_{i})$. The resulting weights will balance $c(\bm{x})$ to the weighted population with density $h(\bm{x})f(\bm{x})$ while $g(\bm{x})$ will only be balanced between the treated and control group. Similar to the ATT, all discussion and methods proposed in Sections 3.2 and 6 can be used for the WATE. However, if the population defined by $h(\bm{x})f(\bm{x})$ has improved overlap compared to the sample population, the set $g(\bm{x})$ required to yield a solution will be smaller than the set required to yield a solution for the ATE. Thus, various $h(\bm{x})$ functions could be used in combination with our proposed constrained optimization problem to improve overlap and yield a solution. Yet, this will modify a large portion of the sample covariate population to the population described by $h(\bm{x})f(\bm{x})$ such that this modified population ideally is of specific interest to the given research question.
Distributional balancing weights
Direct balancing weights in part tend to correspond to estimators with reduced variance because they focus only on balancing linear combinations of the set of basis functions $b(\bm{x}).$ However, if the CATE is not linear in the basis functions, the direct balancing weights estimator is generally biased for the ATE. This has motivated distributional balancing weights are derived to minimize the distance between each of the weighted treated group covariate distributions and the sample covariate distribution (Huling and Mak,, 2024). For distances that are integral probability metrics, the corresponding distributional balancing weights balance all functions in a given class (e.g, weights proposed by Huling and Mak, (2024) balance all functions in a particular Sobolev space). Thus, these weights balance a wider class of functional forms, which is desirable as $\mu_{z}(\bm{X})$ is often unknown. Let $\mathcal{D}$ a distributional distance measure and let $F_{n,z,\bm{w}}(\bm{x})$ be a weighted covariate distribution in treatment group $z$. Then distributional weights commonly take the follow form,
$\bm{w}_{n},=\text{argmin}_{\bm{w}}$ $\>\>\>\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}+\mathcal{D}\{F_{n,1,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}$
where $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ can also be added to minimization objective. While there is always a solution to this optimization problem, when there is a lack of overlap, the corresponding estimator may have inflated variance. To address this, Chen et al., (2024) propose weighting $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}+\mathcal{D}\{F_{n,1,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}$ by a hyperparameter $0\geq\alpha\geq 1$ and $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ by $1-\alpha$; increasing the relative weight of $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ tends to decrease variance but increase bias due to deviation from the target population. When $\alpha=0$, the corresponding weights target the ATO. However, this results in the entire covariate population changing for any $\alpha>0$. Thus, we propose the following distributional weights,
$\bm{w}_{n},=\text{argmin}_{\bm{w}}$ $\>\>\>\mathcal{D}[F_{n,0,\bm{w}}\{c(\bm{x})\},F_{n}\{c(\bm{x})\}]+\mathcal{D}[F_{n,1,\bm{w}}\{c(\bm{x})\},F_{n}\{c(\bm{x})\}]$
$+\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\},$
where only a subset $c(\bm{x})$ of the covariate distribition is balanced to the sample distribution, but the entire covariate distribution is balanced between treatment and control. In addition, $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ could be replaced by $\mathcal{D}[F_{n,0,\bm{w}}\{g(\bm{x})\},F_{n,1,\bm{w}}\{g(\bm{x})\}]$ if desired. Alternatively, Mak and Joseph, (2018) define distributional distances based on a kernel that allows each covariate to impact the distance in different ways; this could be used within a standard distributional balancing weights set-up to weaken the impact of the $g(\bm{x})$ on the distributional distance between the weighted treatment groups and the sample population.
Additional guidance and implementation details for method
Intuition and guidance on selecting the covariates set only balanced between treated and control groups
Selecting the $g(\bm{x})$ set requires researchers to evaluate the trade-offs between estimator variance and bias due to a modified $g(\bm{x})$ target population for their particular application. While these trade-offs are unavoidable when there are positivity violations, it is still desirable to select the $c(\bm{x})$ and $g(\bm{x})$ sets such that the corresponding estimator has minimal error with respect to an estimand of scientific interest, whether this the ATE or the average treatment effect for a modified population. Both subject matter knowledge as well as data driven techniques can be used to identify the $c(\bm{x})$ and $g(\bm{x})$ set; here, we provide further guidance on how a variety of factors and preferences may influence the selection of these two sets.
As mentioned in the previous section, the corresponding estimator for Problem (5) has minimal error with respect to the ATE when $c(\bm{x})$ contains all treatment effect modifiers effect modifiers. When the ATE is of scientific interest, it is therefore desirable to determine the set $g(\bm{x})$ that 1) yields a solution to Problem (5); and 2) minimally modifies the treatment effect. There are a variety of scenarios where set of effect modifiers is likely to be smaller than the set of confounders. Of particular interest are scenarios with high dimensional confounders, which tend to yield overlap issues (D’Amour et al.,, 2021), but where the set of effect modifiers potentially have moderate to low dimension. Subject matter knowledge may indicate which covariates are unlikely to be effect modifiers this which covariates should be in the $g(\bm{x})$. A data-driven procedure could also be used to select $g(\bm{x})$ by estimating how much each covariate influences treatment effect heterogeneity; in fact, we propose an algorithm and corresponding model-based style estimator to do this Section 6.
In other scenarios, either 1) all covariates of interest may be effect modifiers or 2) the size of $g(\bm{x})$ required to ensure a solution to Problem (5) may be large such that some effect modifiers need to be included in the set. From a design-based perspective, $g(\bm{x})$ is the set of covariates that may differ from the original sample population; thus in these cases, it is desirable to choose $g(\bm{x})$ such that the resulting weighted population is still meaningful within the given scientific context. In these cases, $g(\bm{x})$ could be chosen as the set of covariates for which their population is of less interest or relevance to the original research question. However, when there are high dimensional covariates in may be challenging to determine which subset yields a solution to Problem (5) and whose corresponding population may be modified without substantially altering the original research question. As described in Section 3.1, the set $g(\bm{x})$ needs to be predictive of treatment assignment $Z$ in order to substantially relax the positivity assumption. As such, there is some level of predictive capacity, $\eta$, that the covariate functions in $g(\bm{x})$ need to jointly satisfy in order for Problem (5) to yield a solution such that many possible $g(\bm{x})$ sets to select. However, selecting the covariate functions that are individually the most strongly predictive of treatment will tend to result in the smallest of such sets that have predictive capacity $\eta$. While selecting smallest $g(\bm{x})$ set that yields a solution to Problem (5) may not minimize the design-based components of estimator bias, it tends to both 1) simplify the interpretation of resulting the modified population; and 2) be more computationally efficient to operationalize. Thus, in Section 6 we proposed a design-based algorithm for identifying this set.
Furthermore, both model- and design-based perspectives can be used together to select a set $g(\bm{x})$ that best matches the original research objective (e.g., $g(\bm{x})$ contains a mix of covariates that are either unlikely to be effect modifiers or highly predictive of $Z$). Therefore, while a lack of overlap forces researchers to confront trade-offs between estimator variance, statistical bias, and bias due population modification, we propose a variety of intuitive arguments and tools such that researchers can incorporate their subject-matter knowledge and research priorities when navigating these tradeoffs within our proposed optimization procedure.
More details on the proposed model-based estimator
In Section 6 we proposed a model-based estimator. To preserve downstream inference we use the following cross-fit style estimator,
$\hat{\tau}^{CF}_{w^{*}}$ $=\frac{|\mathcal{I}_{1}|}{n}\hat{\tau}^{\mathcal{I}_{1}}+\frac{|\mathcal{I}_{2}|}{n}\hat{\tau}^{\mathcal{I}_{2}},\>\>\>\hat{\tau}^{\mathcal{I}_{1}}=\frac{1}{|\mathcal{I}_{1}|}\sum_{i\in\mathcal{I}_{1}}\{\hat{w}_{i}^{*\mathcal{I}_{2}}Z_{i}Y_{i}-\hat{w}_{i}^{*\mathcal{I}_{2}}(1-Z_{i})Y_{i}\}$ (10)
where $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ each contain half of the data, randomly split, $\hat{w}_{i}^{*\mathcal{I}_{2}}$ are weights computed with $\mathcal{I}_{2}$ and Algorithm 1 with a metric of treatment effect modification, and $\hat{\tau}^{\mathcal{I}_{2}}$ is defined in the same manner by swapping $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$. We implement the absolute value of the one-step doubly robust estimator of treatment effect modification,
$TEM_{j}=\left|\frac{1}{\sum_{i}X_{ij}^{2}}\sum_{i=1}^{n}X_{ij}\left\{(Y_{i}-\hat{\mu}_{Z_{i}}(\bm{X}_{i}))\frac{2Z_{i}-1}{Z_{i}\hat{e}(\bm{X}_{i})+(1-Z_{i})(1-\hat{e}(\bm{X}_{i}))}+\hat{\mu}_{1}(\bm{X}_{i})-\hat{\mu}_{0}(\bm{X}_{i})\right\}\right|,$
proposed by Boileau et al., (2025), as this treatment effect modification measure resulted in estimators with the lowest bias across a majority of the simulation scenarios, compared to the Hines et al., (2023) metric and a linear CATE model (supplementary Figure 2). In this estimator, $\hat{e}(\bm{x}),\hat{\mu}_{z}(\bm{x})$ are propensity score and outcome model estimates. Since propensity score estimates are likely to be extreme in scenarios with a lack of overlap we also explored $TEM^{*}_{j}=\text{abs}[\frac{1}{\sum_{i}X_{ij}^{2}}\sum_{i=1}^{n}X_{ij}\left\{\hat{\mu}_{1}(\bm{X}_{i})-\hat{\mu}_{0}(\bm{X_{i}})\right\}]$ which only requires outcome model estimates. In simulations, estimators corresponding to $TEM_{j}$ tended to perform similarly or better to estimators corresponding to $TEM^{*}_{j}$ (supplementary Figure 3). We use $TEM_{j}$ as the metric in Algorithm 1 as estimators corresponding to $TEM_{j}$ tended to perform similarly or better to estimators corresponding to $TEM^{*}_{j}$ (supplementary Figure 3).
Additional simulation methods
For the data generation propensity score model, $\alpha_{j}=0\pm a$ for $a=0.5/\gamma,1.33/\gamma,2.16/\gamma,3/\gamma$ for $j>0$. Here, $a$ determines the strength of relationship between a given covariate and treatment assignment where there are an equal number of coefficients generated with each of $a=0.5/\gamma,1.33/\gamma,2.16/\gamma,3/\gamma$. As $\gamma$ increases, propensity scores become more extreme and overlap decreases. We set $\mu_{0}(\bm{X})=0.25X_{1}+\cdots+0.25X_{p}$ and $\mu_{1}(\bm{X})=\sum_{i=0}^{p}\beta_{i}X_{i}$. We consider $\theta=0.25,0.75$ as the proportion of covariates are effect modifiers. We set $\beta_{0}=1$ and $\beta_{j}=0.25\pm c$ for $c=0.75\delta,0.6\delta,0.45\delta,0.3\delta,0$. Here, $c$ determines the strength of relationship between a given covariate and $\tau(\bm{X})$; for $(1-\theta)p$ coefficients $\beta_{j}=0$ and there are an approximately equal number of remaining coefficients set as each of $c=0.75\delta,0.6\delta,0.45\delta,0.3\delta$. Then, as $\delta$ increases, treatment effect heterogeneity increases or a given $\theta$. We generated 20 different potential types of covariates for each such scenario with all combinations of $\alpha_{j}$ and $\beta_{j}$ values. See supplementary Table 1 below for the $\gamma,\delta$ values used for each scenario.
$\bm{\gamma}$ $\bm{\delta}$
Low Low-Med Med Low Medium High
20% Treated, p = 20 0.75 1 2 0.50 1 2
40% Treated, p = 20 0.50 0.75 1 0.50 1 2
20% Treated, p = 100 2 4 5 0.25 0.75 1.25
20% Treated, p = 100 1 2 3 0.25 0.75 1.25
Table 1: Data generation overlap and treatment heterogeneity hyperparmeters for each simulation scenario.
Figure 1: Propensity score distribution for the simulation scenarios in Section 7. The scenarios are labeled on the top and right of the panels.
Figure 2: Comparison of estimator MSE with respect to the ATE for our proposed design-based estimator and model-based estimator (with the Boileau et al., (2025), Hines et al., (2023), and linear CATE model coefficient metrics) for both supplementary Algorithm 1 (not-adaptive) and Algorithm 1 (adaptive). The scenarios are labeled on the top and right of the panels.
Figure 3: Comparison of estimator MSE for our model-based estimator with the Boileau et al., (2025) and Hines et al., (2023) metrics calculated with and without estimated propensity scores. The scenarios are labeled on the top and right of the panels.
Figure 4: Estimator mean squared error (MSE) for all 40% treated simulation scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 5: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 20% treated and $p=20$. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 6: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 20% treated and $p=100$. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 7: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 40% treated and $p=20$. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 8: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 40% treated and $p=100$. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Algorithm 1 Identifying an approximately optimal $g(\bm{X})$ set given metric $m$
1:Inputs: Set of functions of covariates ordered according to some metric, $m$, $b_{(1)}(\bm{X}),\ldots,b_{(J)}(\bm{X})$
2:Initialize: $b(\bm{X})^{-}=\emptyset$, $j=0$
3:while $w^{*}$ given $g(\bm{X})=b(\bm{X})^{-}$ does not exist do
4:  Let j = j + 1
5:  Let $b(\bm{X})^{-}=\{b(\bm{X})^{-},b_{(j)}(\bm{X})\}$
6:  Solve for $w^{(}$ in Problem (5) for $g(\bm{X})=b(\bm{X})^{-}$ and $c(\bm{X})=\{b_{(1)}(\bm{X}),\ldots,b_{(J)}(\bm{X})\}\setminus b(\bm{X})^{-}$
7:end while
8:Outputs:
9:   $b(\bm{X})^{-}$ $\triangleright$ Smallest set (given metric $m$) that yields a solution to Problem (5) when $g(\bm{X})=b(\bm{X})^{-}$
10:   $\{w^{*}\}_{i=1}^{n}$ $\triangleright$ Resulting weights from (5) with $g(\bm{X})=b(\bm{X})^{-}$
Figure 9: Distribution of estimated propensity scores for the study on indwelling arterial catheters (IAC) with the data from the MIMIC-III database (Section 8)
Figure 10: Treated and control group standardized mean differences (SMD) for all covariates for the study on indwelling arterial catheters (IAC) with the data from the MIMIC-III database (Section 8).
Figure 11: Treated and sample population standardized mean differences (SMD) for all covariates for the study on indwelling arterial catheters (IAC) with the data from the MIMIC-III database (Section 8).
Figure 12: Control and sample population standardized mean differences (SMD) for all covariates for the study on indwelling arterial catheters (IAC) with the data from the MIMIC-III database (Section 8).
Figure 13: Distribution of estimated probabilities of being in the target population for transporting the health care hotspotting effect to a Midwestern academic health center population (Section 9)
Figure 14: Treated and control group standardized mean differences (SMD) for all covariates for transporting the health care hotspotting effect to a Midwestern academic health center population (Section 9).
Figure 15: Treated and target population standardized mean differences (SMD) for all covariates for transporting the health care hotspotting effect to a Midwestern academic health center population (Section 9).
Figure 16: Control and target population standardized mean differences (SMD) for all covariates for transporting the health care hotspotting effect to a Midwestern academic health center population (Section 9).